### v2

# Model-Free Assessment of Simulator Fidelity via Quantile Curves

Abstract
As generative AI models are increasingly used to simulate real-world systems, quantifying the “sim-to-real” gap is critical. The distributional discrepancy between real and simulated outputs is a random variable driven by the stochastic input scenario. A fundamental challenge is that for any given input, the ground-truth and simulated output distributions are only observable through finite batches of samples, often of heterogeneous sizes. This renders standard predictive inference methods inapplicable, as they seek to quantify uncertainty in observable outputs rather than their underlying population parameters. To address this, we construct confidence sets for these latent parameters and use them to derive a robust proxy for the sim-to-real discrepancy. We then estimate the quantile function of this proxy to provide a comprehensive risk profile of the simulator. Our method is model-agnostic and handles general output spaces, such as categorical survey responses and continuous multi-dimensional sensor data. By rigorously accounting for sampling error, the resulting risk profile supports statistical inference for the real output distribution in a new scenario, the calculation of risk measures like Conditional Value-at-Risk (CVaR), and principled comparisons across simulators. We demonstrate the practical utility of this method by evaluating the alignment of four major LLMs with human populations on the WorldValueBench dataset.
Keywords: Simulation, Quantile function estimation, Human-AI alignment, Conformal inference, Output Uncertainty Quantification
1 Introduction
The adoption of simulators across operations and manufacturing, agent-based modeling in social-science research, user surveys, and education (Zhang et al., 2019; Macal, 2016; Argyle et al., 2023; Aher et al., 2023) has accelerated with recent advances in artificial intelligence (AI). At the platform level, industry ecosystems such as NVIDIA Omniverse and Earth-2 exemplify the push toward high-fidelity, AI-enabled digital twins, and LLMs are increasingly used to build generative agents that emulate human responses (Park et al., 2023; Lu et al., 2025). Collectively, these developments are driving growth in simulation across domains.
Against this backdrop, quantifying the distributional discrepancy between simulated and real-world outcomes is critical. This “sim-to-real” gap is a pervasive challenge in domains ranging from LLMs to computing systems and robotics (Gao et al., 2025; Durmus et al., 2023; Tobin et al., 2017; Peng et al., 2018). Crucially, this gap is not static; it is a function of the scenario (e.g., operating conditions or user prompts). Because a simulator is deployed across a heterogeneous population of scenarios, the sim-to-real gap behaves as a random variable induced by the underlying scenario distribution. Consequently, assessing fidelity requires characterizing the distribution of this random variable, rather than merely reporting a single average error. This is the primary objective of our work.
Existing uncertainty quantification methods that focus on individual outcomes (e.g., conformal prediction) are insufficient for this task, as they do not gauge how well the simulator captures the true output distribution. Conflating outcome validity with distributional fidelity can have serious downstream implications. For instance, in survey research, an LLM-based simulator might generate coherent individual responses while yielding systematically biased population-level aggregates (Westwood, 2025). Recent work has identified distributional alignment as a key objective (Cao et al., 2025), and we address the practical challenge of assessment. We provide theoretical guarantees for sim-to-real discrepancy quantification without making structural assumptions about the simulator, treating it strictly as a black box.
Contributions
We characterize the quantile function of the discrepancy between real and simulated output distributions across scenarios. Because the real-world distribution is only observed through finite, often heterogeneous, batches, the oracle quantile function is not directly computable. We construct per-scenario confidence sets for the latent real-world parameters, use them to form a proxy for the sim-to-real discrepancy, and establish finite-sample guarantees for the resulting quantile curve at any desired level. Moreover, we allow the coverage of the confidence-set to be chosen as a function of the number of sample available. This yields an adaptive procedure that produces tighter quantile curves in data-rich scenarios, while preserving reliability and interpretability when data are scarce.
We illustrate the methodology on WorldValueBench (Zhao et al., 2024), constructed from the World Values Survey (Haerpfer et al., 2020), by generating synthetic response profiles and benchmarking four LLMs against human survey data. Leveraging the large human sample size, we also study tightness and propose a data-driven procedure to construct confidence bands for the population quantile function.
Our main contributions are:
  1. (i)
    Model-free, finite-sample fidelity profiling. A black-box assessment framework with finite-sample, distributional guarantees under heterogeneous batching.
  2. (ii)
    Quantile-based risk summaries. A calibrated discrepancy quantile curve enabling the estimation of tail-risk measures, such as CVaR, and other summaries.
  3. (iii)
    Inference and comparison. Statistical inference for real-world parameters in new scenarios and principled comparisons across simulators.
Related Literature
Figure 1: Simulation Uncertainty Quantification.
The sim-to-real discrepancy has been the focus of the uncertainty quantification (UQ) literature. Figure 1 positions our work within this landscape. Following the taxonomy of Roy and Oberkampf (2011), we distinguish input uncertainty from output uncertainty. Work on input uncertainty (see Chen et al., 2024; Barton et al., 2014; Lam, 2022) typically assumes that the simulator model is known and it produces a faithful representation of the real world when the input follows the correct distribution. The goal is then to characterize how errors in characterizing the input distribution, due to a finite amount of noisy data, propagate to errors in the output. While early work on input uncertainty focused on confidence intervals at a fixed significance level $\alpha$ for functionals of the output, recent work by Chen et al. (2024) constructs Kolmogorov–Smirnov–type uncertainty bands for the CDF of the output discrepancy.
In contrast, modern digital twins rely on complex ML models (e.g., an LLM or a deep neural network), and one often does not have access to the model directly. Moreover, it can be computationally prohibitive to calibrate its parameters exactly. Consequently, we treat the simulator as a black box and focus on directly characterizing the sim-to-real discrepancy of the output—an approach termed output uncertainty quantification by Jeon et al. (2024). A further distinction is that much of the UQ literature is asymptotic, whereas we provide finite-sample guarantees, which is critical when only a limited number of scenarios are available. Finally, while the classical input uncertainty literature is primarily concerned with Monte Carlo noise, which is the variance induced by stochastic sampling conditional on a fixed input model, our work incorporates both variance and bias of the simulator relative to the real world. This is precisely the focus of “sim-to-real” discrepancy.
Recent work in the LLM literature has also begun to study output uncertainty using model-agnostic discrepancy measures. For example, Santurkar et al. (2023a) aggregate the sim-to-real gap into a single scalar summary of overall error or bias, and Huang et al. (2025) bound the discrepancy at a fixed quantile level. These approaches correspond to evaluating particular functionals of the distribution of the sim-to-real discrepancy. More generally, the literature has largely emphasized scalar summaries or fixed-level bounds, leaving open the question of how to obtain a calibrated, distribution-level risk profile across scenarios.
To obtain such distribution-level guarantees in a model-agnostic way, we draw on ideas similar to conformal and predictive inference methods (Vovk et al., 2005; Bates et al., 2021), which provide finite-sample, distribution-free guarantees for black-box predictors. However, standard conformal formulations are typically stated in terms of observable outcomes (e.g., a future label or response) given each input, whereas our target is a scenario-level distributional parameter (e.g., the whole conditional output distribution or its functionals) that is only observed through finite batches, often of heterogeneous sizes. Recent conformal variants that target distributional objects (Snell et al., 2022; Budde et al., 2025) still assume homogeneous data, and do not directly address heterogeneous batching with latent scenario-level targets. We address this setting by providing finite-sample guarantees for quantile-based risk profiles of sim-to-real discrepancy when both real and simulated distributions are observed through heterogeneous batches.
The remainder of the paper proceeds as follows. In Section 2 we present a motivating example, and formally state the problem. In Section 3 we present the main theoretical results, and in Section 4 we discuss an application of our methodology. We also address the tightness of our methodology to the actual quantile curve in Section 5. In Section 6 we conclude by discussing future directions.
2 Framework and Motivating Example
We begin with a motivating use case and then formalize the quantile estimation problem; Appendix D gives additional examples.
Suppose a media research company plans to survey customers on a particular topic with multinomial outcomes (e.g., ‘Agree’, ‘Neutral’, ‘Disagree’). The company wants to estimate customer opinion before committing resources to an expensive population study. The company is considering whether to directly use a digital twin based on pre-trained LLM or to develop its own fine-tuned LLM. In order to make this decision, the company needs to evaluate the distribution of the sim-to-real gap across the range of questions that it is likely to encounter. While pre-trained LLMs can appear to provide coherent predictions, in Appendix E we show an example where several pre-trained LLMs underperform the baseline that randomly picks answers! The methods developed in this work can help with such assessments.
We formalize this setting and the associated theoretical challenges, by considering two sources of uncertainty: variation across scenarios and sampling noise within each scenario. First, we assume that scenarios (questions) are drawn from a population distribution, i.e. $\psi\sim\Psi$, and the past dataset consists of $m$ scenarios $\{\psi_{j}\}_{j=1}^{m}$ drawn i.i.d. from $\Psi$. The real system (human) is characterized by a latent profile $z\in\mathcal{Z}$ distributed according to $\mathcal{P}$ over $\mathcal{Z}$. For each pair $(\psi,z)$, the real system (human) produces a categorical outcome $Y^{\mathrm{gt}}\in\{e_{1},\cdots,e_{d}\}\subset\mathbb{R}^{d},$ with conditional distribution $Q^{\mathrm{gt}}(\cdot\mid z,\psi)$. The simulator (LLM) produces an outcome $Y^{\mathrm{sim}}$ with conditional distribution $Q^{\mathrm{sim}}(\cdot\mid z^{\mathrm{sim}},\psi,r)$ over the same outcome space, where the $z^{\mathrm{sim}}\in\mathcal{Z}_{\mathrm{sim}}$ denotes a synthetic profile drawn i.i.d. from a simulator population distribution $\mathcal{P}^{\mathrm{sim}}$ on $\mathcal{Z}_{\mathrm{sim}}$, and $r$ denotes LLM settings, including prompting strategy, hyperparameters, and other API settings. [^1] We keep $r$ fixed during calibration so that variation in $\hat{q}_{j}$ reflects scenario differences rather than encoding drift or model choice; choosing a different $r$ defines a different simulator. In the running example, $Q^{\mathrm{gt}}(\cdot|z,\psi)=\text{Categorical}(\Pi^{\mathrm{gt}}(z,\psi))$, where $\Pi^{\mathrm{gt}}(\psi,z):=\big(Q^{\mathrm{gt}}(\{1\}\mid z,\psi),\dots,Q^{\mathrm{gt}}(\{d\}\mid z,\psi)\big)\in\mathcal{P}^{d}$ and $\mathcal{P}^{d}$ is the d1d-1-simplex $\{u\in[0,1]^{d}:\sum_{i=1}^{d}u_{i}=1\}$. For any question $\psi$, we can marginalize the population effect, hence denote by $Q^{\mathrm{gt}}(\cdot\mid\psi)=\mathbb{E}_{z\sim\mathcal{P}}[Q^{\mathrm{gt}}(\cdot\mid z,\psi)]$, the conditional distribution of outcome $Y^{\mathrm{gt}}$ given $\psi$.
Let $p(\psi)$ be a population statistic of interest, which is a functional of the conditional distribution $Q^{\mathrm{gt}}(\cdot\mid\psi)$ and lives in a parameter space $\Theta$. In our running example, $\Theta=\mathcal{P}^{d}$, $p(\psi)$ is the mean response of survey respondents, and simulator can be defined similarly under the simulator population $\mathcal{P}^{\mathrm{sim}}$. Further examples of $p(\psi)$ are given in Section 3. To simplify notation, we will write $p(\psi)$ and $q(\psi)$ as $p_{\psi}$ and $q_{\psi}$. Summing up:
$p_{\psi}:=p(\psi)$ $:=\mathbb{E}_{y\sim Q^{\mathrm{gt}}}[y]=\mathbb{E}_{z\sim\mathcal{P}}\big[\Pi^{\mathrm{gt}}(\psi,z)\big]\in\Theta,$
$q_{\psi}:=q(\psi)$ $:=\mathbb{E}_{z\sim\mathcal{P}^{\mathrm{sim}}}\big[\Pi^{\mathrm{sim}}(\psi,z)\big]\in\Theta.$
Second, we only observe finite samples per scenario. For each $j\in[m]$, we are given $n_{j}$ i.i.d. profiles $z_{j,1:n_{j}}\sim\mathcal{P}$, with which we observe $n_{j}$ ground-truth outcomes $y^{\mathrm{gt}}_{j,i}\sim Q^{\mathrm{gt}}(\cdot\mid z_{j,i},\psi_{j})$. Comparably, we generate $k$ simulator outcomes $y^{\mathrm{sim}}_{j,\ell}\sim Q^{\mathrm{sim}}(\cdot\mid z^{\mathrm{sim}},\psi_{j},r)$ using a simulation pool $z^{\mathrm{sim}}_{j,1:k}\sim\mathcal{P}^{\mathrm{sim}}$ with fixed $k$ across $j$ to standardize simulator sampling. For brevity, we will write to be $\hat{p}_{j}$ and $\hat{q}_{j}$ be plug-in estimators of $p_{j}$ and $q_{j}$. In the motivating example these are empirical averages, though the formulation extends to more general plug-in estimators (e.g., empirical distributions or other scenario-level functionals).
Table 1: Illustrative Example of $\mathcal{D}$
Scenario True Distribution ($p_{j}$) $\boldsymbol{n_{j}}$ Real Output ($y^{\mathrm{gt}}$)
11 (0.8, 0.1, 0.1)(0.8,\ 0.1,\ 0.1) 1010 (1, 1, 2, 3, 1, 2, 1, 1, 1, 1)(1,\ 1,\ 2,\ 3,\ 1,\ 2,\ 1,\ 1,\ 1,\ 1)
22 (0.3, 0.4, 0.3)(0.3,\ 0.4,\ 0.3) 44 (3, 1, 2, 2)(3,\ 1,\ 2,\ 2)
$\vdots$ $\cdots$ $\vdots$ $\cdots$
$m$ (0.05, 0.05, 0.9)(0.05,\ 0.05,\ 0.9) 88 (1, 3, 3, 3, 3, 2, 2, 3)(1,\ 3,\ 3,\ 3,\ 3,\ 2,\ 2,\ 3)
The dataset is $\mathcal{D}=\big\{(\psi_{j},\hat{p}_{j},\hat{q}_{j},n_{j},k)\big\}_{j=1}^{m}$. For each scenario $j$, let $\mathcal{D}^{\mathrm{gt}}_{j}=\{(z_{ji},y^{\mathrm{gt}}_{ji})\}_{i=1}^{n_{j}}$ denote the human-side data used to construct $\hat{p}_{j}$, and $\mathcal{D}^{\mathrm{sim}}_{j}=\{(z^{\mathrm{sim}}_{jl},y^{\mathrm{sim}}_{jl})\}_{l=1}^{k}$ denote the simulator outputs used to construct $\hat{q}_{j}$. For concreteness, Table 1 provides an illustrative example of $\mathcal{D}$. Aligning with our motivating survey example, $p_{j}$ is a multinomial parameter and the human sample sizes $\{n_{j}\}$ may vary across scenarios, so the realized outcomes $\{y^{\mathrm{gt}}_{ji}\}$ differ both in content and in length. We record outcomes as category indices in $[d]$; equivalently, one may use one-hot encoding $e_{y}\in\mathbb{R}^{d}$. For instance, when $j=2$ the true distribution is $p_{2}=(0.3,0.4,0.3)$ and a sample of size $n_{2}=4$ yields $\hat{p}_{2}=(0.25,0.5,0.25)$.
The discrepancy for a specific question $\psi$ between simulated and real output distributions is defined as
$\Delta_{\psi}:=L(p_{\psi},\hat{q}_{\psi}),$
where $L:\Theta\times\Theta\to[0,\infty)$ is a user-chosen discrepancy function. We define the discrepancy in terms of $\hat{q}_{\psi}$ to reflect a fixed simulator query budget; when simulator sampling is effectively unlimited, one may replace $\hat{q}_{\psi}$ with $q_{\psi}$ to study the intrinsic discrepancy $L(p_{\psi},q_{\psi})$. We discuss this further in Section 3. Our method also works with any function $L$ and allows practitioners to choose one that suits their application, e.g. Kullback–Leibler (KL) divergence for categorical outputs or a Wasserstein distance for nonparametric empirical distributions.
Now, we can state our task:
Construct a function $\hat{V}(\cdot,\mathcal{D}):(0,1)\to\mathbb{R}$ such that, for a new $\psi\sim\Psi$ and all $\alpha\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta_{\psi}\leq\hat{V}(\alpha,\mathcal{D})\big|\mathcal{D}\big)\geq\alpha-\varepsilon_{m},$
holds with high probability over the draw of $\mathcal{D}$ and $\lim_{m\rightarrow\infty}\varepsilon_{m}=0$.
Note that $\hat{V}$ is indexed by both the prescribed level $\alpha$ and the dataset $\mathcal{D}$. Establishing our desired result requires controlling two sources of uncertainty: (i) scenario uncertainty, since we observe only finitely many scenarios $\{\psi_{j}\}_{j=1}^{m}\sim\Psi$, and (ii) finite-sample uncertainty within each scenario, since $p_{j}$ and $q_{j}$ are only observed through finite samples that produce $\hat{p}_{j}$ and $\hat{q}_{j}$. Our construction of $\hat{V}(\alpha,\mathcal{D})$ explicitly accounts for both, and the main theorem will yield non-vacuous, finite-sample coverage guarantees.
3 Theoretical Result
In this section, we construct $\hat{V}$, establish its theoretical guarantees, and conclude by discussing the benefits of distribution-level fidelity guarantees.
3.1 Methodology
Our dataset $\mathcal{D}$ consists of $m$ scenarios; the true parameter value (resp. finite sample estimate) for scenario $j$ is $p_{j}:=p(\psi_{j})$ (resp. $\hat{p}_{j}$); the simulator parameter value (resp. finite sample estimate) for scenario $j$ is $q_{j}:=q(\psi_{j})$ (resp. $\hat{q}_{j}$).
Our procedure has two steps:
  1. 1.
    For each $j$, define a $\gamma_{j}$-confidence set $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$, i.e.
    $\mathbb{P}(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma_{j}\right))\geq\gamma_{j},$ (3.1)
    and the pseudo-discrepancy
    $\hat{\Delta}_{j}:=\sup_{u\in\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma_{j}\right)}L\big(u,\hat{q}_{j}\big).$ (3.2)
  2. 2.
    For $\alpha\in(0,1)$, let
    $\hat{V}_{m}\!\left(\alpha\right):=\text{the $\alpha$-quantile of}\{\hat{\Delta}_{j}\}_{j=1}^{m}.$ (3.3)
We will show that $\hat{V}_{m}\!\left(\alpha\right)$ is a calibrated quantile curve for the discrepancy distribution, up to explicit finite-sample corrections. Note that the key design choice in our method is the confidence set $\mathcal{C}_{j}\!\left(\cdot\right)$ that quantifies the uncertainty of $\hat{p}_{j}$. This uncertainty set is the key device that links per-scenario sampling error to distribution-level calibration. Next, we show how to construct these sets using off-the-shelf concentration bounds, starting with the multinomial setting.
Example 3.1 (Multinomial Confidence Set).
Suppose the outcomes are multinomial with parameter $p_{j}\in\Theta=\mathcal{P}^{d}:=\{u\in[0,1]^{d}:\sum_{i=1}^{d}u_{i}=1\}$, i.e., $Y_{j,i}\mid\psi_{j}\sim\mathrm{Categorical}(p_{j})$ and $\hat{p}_{j}$ is the empirical average vector from $n_{j}$ samples. Let $\mathrm{KL}(\cdot\|\cdot)$ denote the KL divergence. Then
$\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma_{j}\right):=\Big\{u\in\mathcal{P}^{d}:\mathrm{KL}(\hat{p}_{j}\|u)\leq\frac{d-1}{n_{j}}\ln(\frac{2(d-1)}{1-\gamma_{j}})\Big\}.$
is a $\gamma_{j}$-confidence set for $p_{j}$.
This result is a variant of Chernoff-Hoeffding Bound, see Lemma A.5 (Mardia et al., 2019).
Example 3.2 (Bounded Outcomes).
Suppose $Y_{j,i}\in[a,b]$ are i.i.d. with population mean $p_{j}:=\mathbb{E}[Y_{j,i}]$ for scenario $j$, and $\hat{p}_{j}:=\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}Y_{j,i}$ is the sample mean. Then
$\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma_{j}\right):=\Big\{u\in[a,b]:|u-\hat{p}_{j}|\leq(b-a)\sqrt{\tfrac{\log(2/(1-\gamma_{j}))}{2n_{j}}}\Big\},$
is a $\gamma_{j}$-confidence set for $p_{j}$.
Example 3.3 (Nonparametric $W_{1}$ Confidence Set).
Suppose the true outcome distribution for scenario $j$ is $P_{j}$ and samples are drawn as $Y_{j,1:n_{j}}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}P_{j}$. Assume $P_{j}$ is $\sigma$-sub-Gaussian. Let $\widehat{P}_{j}$ denote the empirical distribution from $n_{j}$ samples. Then
$\mathcal{C}_{j}^{W_{1}}(\widehat{P}_{j},\gamma_{j}):=\Big\{Q:\ W_{1}(\widehat{P}_{j},Q)\leq r_{j}(n_{j},\gamma_{j},\sigma)\Big\},$
where
$r_{j}=\frac{512\sigma}{\sqrt{n_{j}}}+\sigma\sqrt{\frac{256e}{n_{j}}\log\frac{1}{1-\gamma_{j}}},$
is a $\gamma_{j}$-confidence set for the true distribution $P_{j}$.
The form of $\mathcal{C}_{j}$ is justified by the concentration inequality proved in L.A. and Bhat (2022).
Next, we show that $\hat{\Delta}_{j}$ can be efficiently computed for each scenario $j$.
  1. (a)
    KL-Divergence: For $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})=\{u\in\mathcal{P}^{d}:\ \mathrm{KL}(\hat{p}_{j}\|u)\leq r_{j}\}$ and loss $L(u,\hat{q}_{j})=\mathrm{KL}(u\|\hat{q}_{j})$, the pseudo-discrepancy
    $\hat{\Delta}_{j}=\sup_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})}\mathrm{KL}\big(u\|\hat{q}_{j}\big).$
    The maximizer lies on the boundary $\mathrm{KL}(\hat{p}_{j}\|u)=r_{j}$ and can be computed via standard maximization.
  2. (b)
    Wasserstein–1: For $\mathcal{C}^{W_{1}}_{j}(\widehat{P}_{j},\gamma_{j})=\{u:W_{1}(u,\widehat{P}_{j})\leq r_{j}\}$ and loss $L(u,\widehat{Q}_{j})=W_{1}(u,\widehat{Q}_{j})$, triangle inequality implies that
    $\hat{\Delta}_{j}=\sup_{u\in\mathcal{C}^{W_{1}}_{j}(\widehat{P}_{j},\gamma_{j})}W_{1}(u,\widehat{Q}_{j})\leq W_{1}(\widehat{P}_{j},\widehat{Q}_{j})+r_{j}.$
3.2 Calibrated Quantile Curve Theory
Our theoretical guarantee relies on the following two assumptions.
Assumption 3.1 (Independent data).
The scenarios $\psi_{j}\sim\Psi$ are i.i.d. , $\mathcal{D}^{\mathrm{gt}}_{j}\perp\!\!\!\perp\mathcal{D}^{sim}_{j}$ conditioned on $\psi_{j}$, and the new $(\psi,\mathcal{D}^{sim})$ is independent of $\mathcal{D}$. In addition, the human sample sizes $\{n_{j}\}_{j=1}^{m}$ are treated as fixed (deterministic) design quantities.
Assumption 3.2 (Regular Discrepancy).
The discrepancy $L:\Theta\times\Theta\to[0,\infty)$ is jointly continuous on $\Theta\times\Theta$ and satisfies $L\!\left(u,u\right)=0$ for all $u\in\Theta$.
Before stating our main result, we consider the special case $\gamma_{j}\equiv\tfrac{1}{2}$. Then, in the asymptotic regime $m\to\infty$, our guarantee (3.6) below roughly implies the following: for any target quantile level $\tau\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}\leq\hat{V}_{m}\Big(\tfrac{1+\tau}{2}\Big)\Big|\mathcal{D}\Big)\geq\tau.$ (3.4)
In words, the $\tau$-quantile of the discrepancy $\Delta_{\psi}$ is approximated by the empirical pseudo-quantile curve at an inflated index $(1+\tau)/2$. This inflation induces an inherent conservativeness: for $\tau\approx 0$, the index $(1+\tau)/2$ is at least the median of the pseudo-discrepancy distribution, leading to substantial inflation in the lower tail.
In general, the $\gamma_{j}$’s can be different, and we obtain the following result:
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}\leq\hat{V}_{m}\big(\bar{\gamma}_{m}\tau+(1-\bar{\gamma}_{m})\big)\Big|\mathcal{D}\Big)\geq\tau,$ (3.5)
where $\bar{\gamma}_{m}=\tfrac{1}{m}\sum_{j=1}^{m}\gamma_{j}$. Now, to achieve $\tau$-level coverage we evaluate the empirical pseudo-quantile at an index that is affine in $\tau$ with slope $\bar{\gamma}_{m}$. One would ideally want $\bar{\gamma}_{m}=1$, or equivalently $\gamma_{j}\equiv 1$, to control the index inflation. However, increasing $\gamma_{j}$ typically enlarges the set $\mathcal{C}_{j}$, and thus, increases the values of the pseudo-discrepancies $\hat{\Delta}_{j}$ and shift the curve $\hat{V}_{m}$ upward. Thus, one needs to balance index adjustment with increased conservativeness of the pseudo-quantile curve itself.
This trade-off motivates choosing $\gamma_{j}$ adaptively as a function of the human sample size $n_{j}$. When $n_{j}$ is large, $p_{j}$ is typically more concentrated around $\hat{p}_{j}$; therefore, $\gamma_{j}$ can be set high without significantly enlarging $\mathcal{C}_{j}$. Therefore, we set $\gamma_{j}=g(n_{j})$, for an increasing function $g:\mathbb{R}_{+}\to[0,1]$. See Section 4 for numerical evidence illustrating the conservativeness of fixed $\gamma$ and the benefits of adaptive coverage. We are now ready to rigorously state the main theorem.
Theorem 3.1.
Suppose Assumptions 3.13.2 hold and $\gamma_{j}\in(0,1)$ for each $j\in[m]$. Let $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$ denote any $\gamma_{j}$-confidence set (see (3.1)), $\hat{\Delta}_{j}$ denote the pseudo-discrepancy (see (3.2)), and let $\hat{V}_{m}$ denote the quantile function of the pseudo-discrepancy (see (3.3)). For $\alpha\in(0,1)$ let $\alpha]_{m}:=m^{-1}\lceil m\alpha\rceil\in\Big[\alpha,\alpha+\frac{1}{m}\Big]$ and set
$\alpha_{\mathrm{eff}}(\alpha):=\Big[\big(\bar{\gamma}_{m}-\varepsilon_{\mathrm{al}}(\alpha,\delta)\big)[\alpha]_{m}-b_{m}(\delta)\sqrt{[\alpha]_{m}}\Big]_{+},$
where $\bar{\gamma}_{m}:=\frac{1}{m}\sum_{j=1}^{m}\gamma_{j}$, $\varepsilon_{\mathrm{al}}(\alpha,\delta)=\sqrt{\frac{\log(3m/\delta)}{2m[\alpha]_{m}}}$, and $b_{m}(\delta)=\sqrt{\frac{1}{2m}\log\frac{3m}{\delta}}$. Then, with probability at least $1-\delta$ over $\mathcal{D}$, simultaneously for all $\alpha\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}^{(k)}\leq\hat{V}_{m}\big(1-\alpha_{\mathrm{eff}}(\alpha)\big)\Big|\mathcal{D}\Big)\geq 1-\alpha-\varepsilon_{m}(\delta),$ (3.6)
where $\varepsilon_{m}(\delta):=\sqrt{\frac{\log(6/\delta)}{2m}}+\tfrac{1}{m}$. In particular, the remainder terms $\varepsilon_{m},\varepsilon_{\mathrm{al}},$ and $b_{m}\to 0$ as $m\to\infty.$
Unlike standard conformal inference, which targets uncertainty quantification for future observable outcomes (e.g. responses), Theorem 3.1 provides uncertainty quantification for the latent, scenario-specific population parameter that governs outcome generation (e.g. conditional mean response). In other words, the guarantee operates at the level of underlying distributions, in contrast to single draws. We next provide a proof sketch.
Proof Sketch:
Suppose we have access to the true $p_{j}$ for each scenario $j$. Then $\{\Delta_{j}\}_{j=1}^{m}$ are i.i.d. , and the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality yields a uniform deviation bound for the empirical CDF of the discrepancy, which can be translated into upper bounds on its quantile function. In our setting, however, (i) $p_{j}$ is latent and (ii) each scenario is observed with a heterogeneous sample size $n_{j}$.
Our goal is to use empirical quantiles of the observable proxy $\{\hat{\Delta}_{j}\}$ to approximate the quantile function of the unobservable discrepancies $\{\Delta_{j}\}$. For simplicity, we fix $\gamma_{j}\equiv\gamma$. The key steps are as follows:
  • Let $t_{\alpha}$ denote the population $(1-\alpha)$-quantile of $\Delta_{j}$, and define the corresponding hardest scenarios $S_{\alpha}:=\{j:\Delta_{j}\geq t_{\alpha}\}$. By definition, $\tfrac{|S_{\alpha}|}{m}\approx\alpha$.
  • By construction, $\hat{\Delta}_{j}\geq\Delta_{j}$ holds with probability $\gamma$ for each $j$. Consequently, among indices in $S_{\alpha}$, at least a $\gamma$-fraction satisfy $\hat{\Delta}_{j}\geq t_{\alpha}$.
  • Hence, at least a $\gamma\alpha$-fraction of the pseudo-discrepancies exceed $t_{\alpha}$. In other words, the $(1-\alpha)$ quantile of $\Delta_{j}$ is dominated by the $(1-\gamma\alpha)$ quantile of $\hat{\Delta}_{j}$.
Therefore, for any quantile level $\tau:=1-\alpha$, taking the $(\gamma\tau+(1-\gamma))$-quantile of the pseudo-discrepancies yields an upper bound on the $\tau$-quantile of the oracle discrepancies.
This is why the quantile index must shift: to upper bound the oracle $\tau$-quantile of $\Delta$, we need to use a more conservative pseudo-quantile level so that the resulting pseudo-threshold is high enough to cover the target quantile despite the miscoverage. This index adjustment is the first source of conservativeness; the second source is the DKW deviation that relates empirical and population CDFs over finitely many scenarios. Combining these two ingredients yields the theorem. The complete proof is in Appendix B.1. ∎
Remark 1.
A subtle modeling choice is that Theorem 3.1 characterizes the distribution of $\Delta^{(k)}=L(p,\hat{q}^{(k)})$, where $\hat{q}^{(k)}$ is estimated using $k$ simulator samples, instead of the actual simulator parameter $q$. Conceptually, $\Delta^{(k)}$ is the correct target when one wants to characterize performance for a fixed query budget $k$. In contrast, $\Delta$ is more natural when simulator queries are cheap, and the goal also includes quantifying the simulator’s inherent bias. For brevity, we will, henceforth, ignore the $k$ superscript in following discussions. A proof to the latter case is presented in Corollary B.2, see Appendix B for details.
3.3 Benefits of Distribution-Level Characterization
We now turn to the benefits of obtaining a distributional-level characterization of sim-to-real discrepancy. For illustrative purposes we set $\bar{\gamma}_{m}=\tfrac{1}{2}$. With this calibrated quantile, for an unseen scenario $\psi$ we can derive a confidence set for $p_{\psi}$ given a simulated $\hat{q}_{\psi}$. For a target confidence level $\bar{\alpha}\in(0,1)$, define
$S_{\bar{\alpha}}:=\big\{u\in\Theta:L(u,\hat{q}_{\psi})\leq\tau_{\bar{\alpha}}\big\},\quad\tau_{\bar{\alpha}}:=\hat{V}_{m}\!\left(1-\tfrac{\bar{\alpha}}{2}\right).$
By Theorem 3.1, $p_{\psi}\in S_{\bar{\alpha}}$ with probability at least $1-\bar{\alpha}$ up to an $o_{m}(1)$ remainder.
In addition, the calibrated quantile curve can be compressed to provide summary statistics. Returning to the motivating example, a media research firm can use a calibrated AUC to assess average simulator bias, while a pollster or risk-averse product team may track a calibrated $\mathrm{CVaR}_{\alpha}$ to bound rare but consequential misreads of public sentiment.
Concretely, define the index–adjusted (calibrated) curve
$\hat{V}_{m}^{\mathrm{cal}}(\tau):=\hat{V}_{m}\Big(\tfrac{1+\tau}{2}\Big),\qquad\tau\in[0,1],$
which matches the coverage guarantee in Equation 3.4 in asymptotic regime. We then summarize overall average sim-to-real discrepancy via the calibrated AUC
$\mathrm{AUC}_{\mathrm{cal}}:=\int_{0}^{1}\hat{V}_{m}^{\mathrm{cal}}(\tau)\,d\tau,$
and tail risk via a calibrated right-tail CVaR: for $\alpha\in(0,1)$,
$\mathrm{CVaR}^{\mathrm{cal}}_{\alpha}:=\frac{1}{\alpha}\int_{1-\alpha}^{1}\hat{V}_{m}^{\mathrm{cal}}(u)\,du.$
In words, $\mathrm{AUC}_{\mathrm{cal}}$ aggregates the entire calibrated curve into a single average-bias summary, while $\mathrm{CVaR}^{\mathrm{cal}}_{\alpha}$ averages over the worst $\alpha$ fraction of scenarios under the same index-adjusted quantile curve.
4 Application: LLM Fidelity Profiling
4.1 Dataset and Methodology
We evaluate our procedure on the WorldValueBench dataset, curated by Zhao et al. (2024) and constructed from the World Values Survey (Haerpfer et al., 2020). The dataset comprises survey questions asked across 64 countries, covering 12 thematic categories (e.g., social values, security, migration). Individual-level covariates are available for each respondent, including gender, age, migration status, education, marital status, and related demographics.
Figure 2: Example of World Value Question. Source: Haerpfer et al. (2020).
Preprocessing.
After data cleaning, we retain 235 distinct questions and responses from 96,220 individuals. Each question offers a categorical set of possible answers, as illustrated in Figure 2. To place heterogeneous answer sets on a common scale, we map each question’s categories to the interval [1,1][-1,1], producing a bounded real-valued outcome for every scenario and respecting the natural ordering of response options when available (e.g., disagree–neutral–agree). Individual-level covariates are then used to construct synthetic profiles and corresponding prompts for simulators. Additional details on preprocessing and dataset characteristics are provided in Appendix C.
Methodology
We generate synthetic responses and estimate $\{\hat{q}_{j}\}_{j=1}^{235}$ for four LLMs: GPT-4o (gpt-4o), GPT-5 mini (gpt-5-mini), Llama 3.3 70B (Llama-3.3-70B-Instruct-Turbo), and Qwen 3 235B (Qwen3 235B A22B Thinking 2507 FP8). As a benchmark, we also construct a uniform baseline: for each question, this generator samples an answer uniformly at random from the available choices for each individual. We simulate $200$ LLM samples per question, and use these simulated responses to estimate $\hat{q}_{j}$. For humans, we subsample 500 respondents per question to calculate $\{\hat{p}_{j}\}_{j=1}^{235}$ and construct the corresponding confidence sets $\mathcal{C}_{j}$. Due to nonresponses, the effective sample size $n_{j}$ varies across questions (typically 450–500).
Since the outcome is bounded in [1,1][-1,1], we construct per-scenario confidence sets $\mathcal{C}_{j}$ using Example 3.2. We set $\gamma_{j}=1-n_{j}^{-1/3},$ which is increasing in sample size $n_{j}$ [^2] As discussed in Section 3, the coverage level $\gamma_{j}=g(n_{j})$ should increase with the sample size $n_{j}$. We therefore adopt a generic form $\gamma_{j}=1-n_{j}^{-\beta},$ and conduct a sensitivity analysis on $\beta$ value in Appendix C.4., and yields an average coverage level of $\bar{\gamma}_{m}\approx 0.87$. The loss function $L(p,q)=(p-q)^{2}$. Next, we use the procedure described in Section 3 (see (3.1)-(3.3)) to obtain a fidelity profile for each candidate LLM, and plot the calibrated curve using the asymptotic index adjustment in (3.5). For completeness, we also apply the same workflow to a Bernoulli setting (EEDI) and a multinomial setting (OpinionQA); see Appendix E.
4.2 Fidelity Profiling
Figure 3: Calibrated $\hat{V}(\tau)$ across LLMs.
The calibrated quantile functions for the four different LLMs are plotted in Figure 3. Lower, flatter curves indicate uniformly small discrepancies, while elbows reveal rare but severe misses. Our results suggest that all LLM-simulators outperform the uniform benchmark on over $80\%$ of questions, but fail to capture the tail scenarios. GPT-4o is the lowest across all quantiles, indicating the most reliable alignment, with GPT-5-mini close behind yet failing to capture some outlier questions, and Llama 3.3 and QWen-3 trail behind.
Figure 4: Robustness check of simulator performances.
We also provide a robustness check on whether the dominance relation is an anomaly of the particular choice for the sample size $n_{j}$. Figure 4 shows performance of the LLMs for $n_{j}=\{50,500,5000,10000\}$. The results indicate that the dominance relation between LLMs remains the same.
5 Tightness Analysis of Calibrated Quantiles
A natural concern with our approach is the tightness of our calibrated quantile curve. Since our construction relies on off-the-shelf concentration bounds and worst-case pseudo-discrepancies, the resulting calibrated quantiles could, in principle, be conservative to the point of being vacuous. To assess this empirically, we conduct a tightness study in a setting that leverages a feature of WorldValueBench.
We leverage the WorldValueBench benchmark, which contains up to 96,22096{,}220 respondents per question. We treat the empirical distribution $\hat{p}_{j}$ computed from all available respondents as a proxy for the true human response distribution $p_{j}$. This enables computation of an “oracle” discrepancy $\Delta_{j}^{\star}:=L(p_{j},\hat{q}_{j}),$ and hence an oracle quantile curve across questions. We fix the simulator budget at $k=200$ and, for each question $j$, subsample $n_{j}\in\{50,500,1000,5000\}$ respondents to construct the confidence sets and pseudo-discrepancies $\{\hat{\Delta}_{j}\}_{j=1}^{m}$. We then compare the resulting calibrated curves against the oracle quantile curve.
Figure 5: Tightness analysis for varying $n_{j}$ under GPT-4o.
The results plotted in Figure 5 indicate that the calibrated envelope is substantially conservative when $n_{j}=50$, but becomes tighter when $n_{j}\geq 500$. This behavior is consistent with our construction: larger $n_{j}$ yields more concentrated estimates $\hat{p}_{j}$ and smaller confidence sets $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$, which reduces the inflation induced by the supremum defining $\hat{\Delta}_{j}$.
The results in Figure 5 correspond to $\gamma_{j}=1-n_{j}^{-\frac{1}{3}}$. In order to isolate the impact of batch-size-specific $\gamma_{j}$, as we discussed in Section 3, we repeat the same experiment by setting $\gamma_{j}\equiv\tfrac{1}{2}$ for all questions. The corresponding results are plotted in Figure 6. The plots here exhibit a persistent vertical separation between the oracle quantile curve and the calibrated curves, even for large $n_{j}$. This phenomenon is explained by the fact that here $\bar{\gamma}_{m}=\tfrac{1}{2}$ (independent of $n_{j}$) and the asymptotic calibration evaluates the pseudo-quantile at $\hat{V}_{m}\big(0.5\tau+0.5\big)$. Consequently, at lower quantile levels ($\tau\approx 0$), the our procedure approximates the quantile function by at least the median, inducing a non-vanishing conservativeness that cannot be eliminated by increasing $n_{j}$. This underscores the importance of allowing scenario-dependent coverage levels.
Figure 6: Tightness analysis for varying $n_{j}$ under GPT-4o with fixed $\gamma_{j}=\tfrac{1}{2}$.
While WorldValueBench provides a large pool of respondents that facilitates an oracle tightness check, most applications do not enjoy such extensive ground-truth data. We next develop a complementary tightness assessment: a confidence band that sandwiches the oracle quantile curve.
Theorem 5.1 (Informal Version).
Suppose the setup of Theorem 3.1 and Assumptions 3.13.2 hold. Let $\{\gamma_{L,j},\gamma_{U,j}\}_{j=1}^{m}\subset(0,1]$ be coverage levels and define the lower and upper pseudo-discrepancies
$\hat{\Delta}_{j}^{-}:=\inf_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{L,j})}L(u,\hat{q}_{j}),\ \hat{\Delta}_{j}^{+}:=\sup_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{U,j})}L(u,\hat{q}_{j}),$
where $\mathcal{C}_{j}(\hat{p}_{j},\gamma)\subset\Theta$ are $\gamma$-confidence sets. Let $\hat{V}_{m}^{-}(\cdot)$ (resp. $\hat{V}^{+}(\alpha)$) denote the empirical quantile function of $\{\hat{\Delta}_{j}^{-}\}_{j=1}^{m}$ ( resp. $\{\hat{\Delta}_{j}^{+}\}_{j=1}^{m}$). Let $V:[0,1]\to\mathbb{R}$ denote the quantile function of $\Delta(\psi)$. Then, as $m\to\infty$, for any fixed $\tau\in(0,1)$,
$\hat{V}_{m}^{-}\big(\bar{\gamma}_{L,m}\tau\big)\lesssim V(\tau)\lesssim\hat{V}_{m}^{+}\big(\bar{\gamma}_{U,m}\tau+(1-\bar{\gamma}_{U,m})\big),$ (5.1)
where $\bar{\gamma}_{L,m}=\tfrac{1}{m}\sum_{j=1}^{m}\gamma_{L,j}$, and $\bar{\gamma}_{U,m}=\tfrac{1}{m}\sum_{j=1}^{m}\gamma_{U,j}$.
Figure 7: Confidence Bands of GPT-4o and Llama.
The formal version of this result is presented in Theorem B.1 and its proof is deferred to Appendix B.4. The takeaway is that Theorems 3.1 and B.1 provide complementary tools for assessing simulator fidelity at the distributional level. Theorem 3.1 provides a one-sided, calibrated upper envelope for the population quantile curve, while Theorem B.1 augments this with a lower envelope, producing a confidence band that sandwiches the unknown quantile function $V(\tau)$. Figure 7 illustrates these bands for two subsample sizes $n_{j}\in\{1000,5000\}$ and compares GPT-4o against Llama 3.3. Across both panels, the GPT-4o band lies systematically below the Llama 3.3 band over most quantile levels, with the separation becoming most pronounced in the upper tail. This pattern indicates that GPT-4o achieves systematically smaller calibrated discrepancy across scenarios and exhibits fewer extreme failures, and the conclusion is stable as $n_{j}$ increases. Importantly, the comparison is fully data-driven: unlike oracle evaluations that rely on near-population human samples, the band construction does not require access to ground-truth discrepancies and can be carried out using only the observed heterogeneous batches.
6 Discussion
We develop a model-agnostic framework for profiling simulator fidelity through the quantile function of the sim-to-real discrepancy. Our approach makes no parametric assumptions about either the simulator or the real system, rigorously accounts for finite and heterogeneous sampling across scenarios, and yields calibrated, finite-sample guarantees at any target quantile level. The resulting fidelity profile supports a range of downstream summaries, from average performance to tail-risk measures that isolate rare but consequential failures. Finally, we demonstrate the practical value of the framework on real survey data, evaluating multiple LLM-based simulators and illustrating systematic, distribution-level differences in their alignment with human populations.
Despite its broad applicability, the method we introduce leaves several natural extensions. First, our proof relies on DKW-type concentration – often conservative for small $m$ – and a grid-uniform step that further loosens constants; tightening these bounds is an immediate priority. Second, many applications involve temporally dependent, dynamic simulation processes; extending our static framework to dynamic settings would broaden applicability. Third, our analysis assumes i.i.d. scenarios, whereas covariate shift or endogenous sampling may invalidate marginal guarantees; addressing such distribution shifts is an important avenue for future work.
Acknowledgement
Kaizheng Wang’s research is supported by NSF grants DMS-2210907 and DMS-2515679 and a Data Science Institute seed grant SF-181 at Columbia University.
Appendix A Technical Lemmas
Lemma A.1.
(Dvoretzky-Kiefer-Wolfowitz (DKW) Inequality via Massart (1990))
Let $X_{1},X_{2},\dots,X_{n}$ be i.i.d. real-valued random variables with cumulative distribution function (CDF) $F^{*}$, and let $\hat{F}_{n}$ be the empirical distribution function defined by
$\hat{F}_{n}(x):=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{X_{i}\leq x\}.$
Then, for any $\varepsilon>0$,
$\mathbb{P}\left(\sup_{x\in\mathbb{R}}\left|\hat{F}_{n}(x)-F^{*}(x)\right|>\varepsilon\right)\leq 2e^{-2n\varepsilon^{2}}.$
Equivalently, for any confidence level $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sup_{x\in\mathbb{R}}\left|\hat{F}_{n}(x)-F^{*}(x)\right|\leq\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)}.$
Lemma A.2.
[Hoeffding (additive) via Hoeffding (1963)]
Let $Z_{1},\dots,Z_{n}\in[0,1]$ be independent, $T=\sum_{i=1}^{n}Z_{i}$, and $\mu=\mathbb{E}[T]$. For any $t\in[0,\mu]$,
$\mathbb{P}\big(T\leq\mu-t\big)\leq\exp\left(-\frac{2t^{2}}{\sum_{i=1}^{n}(1-0)^{2}}\right)=\exp\left(-\,\frac{2t^{2}}{n}\right).$
Lemma A.3.
[Chernoff–Hoeffding for one-parameter exponential family]
Let $X_{1},\dots,X_{n}$ be i.i.d. with density (or mass) in the one-parameter canonical exponential family
$p_{\theta}(x)\;=\;\exp\{\theta T(x)-A(\theta)\}\,h(x),\qquad\theta\in\Theta,$
where $T(x)$ is the sufficient statistic, $A(\theta)$ is the log-partition function (convex, differentiable on $\Theta$), and the mean map is $\mu(\theta):=\mathbb{E}_{\theta}[T(X)]=A^{\prime}(\theta)$. Assume $\Theta$ is an open interval and all quantities below are finite.
Define the empirical mean of the sufficient statistic
$\bar{T}_{n}\;=\;\frac{1}{n}\sum_{i=1}^{n}T(X_{i}).$
For each $t$ in the range of $\bar{T}_{n}$ let $\theta_{t}$ be the (unique) canonical parameter satisfying $\mu(\theta_{t})=t$. Then for any $\varepsilon>0$,
$\mathbb{P}\bigl(D(p_{\theta_{\bar{T}_{n}}}\|p_{\theta})>\varepsilon\bigr)\leq 2e^{-n\varepsilon}$
Proof:
Define the shifted log-MGF under $p_{\theta}$,
$\psi_{\theta}(\lambda):=\log\mathbb{E}_{\theta}\big[e^{\lambda T(X)}\big]=A(\theta+\lambda)-A(\theta),$
where the displayed equality follows from the exponential-family form (for $\lambda$ in the domain where the expectation is finite). For an attainable mean $m$ denote $\theta_{m}$ as the unique solution of $\mu(\theta_{m})=m$.
By adopting the one-sided Chernoff bound, for any real $\lambda$ such that expectations exist and any $m\in\mathbb{R}$,
$\mathbb{P}_{\theta}(\bar{T}_{n}\geq m)$ $=\mathbb{P}\big(e^{\lambda n\bar{T}_{n}}\geq e^{\lambda nm}\big)$
$\leq e^{-\lambda nm}\,\mathbb{E}_{\theta}\!\big[e^{\lambda n\bar{T}_{n}}\big]$
$=\exp\big(-n(\lambda m-\psi_{\theta}(\lambda))\big).$
Optimizing over $\lambda$ gives the Chernoff bound
$\mathbb{P}_{\theta}(\bar{T}_{n}\geq m)\leq\exp\!\big(-n\psi_{\theta}^{*}(m)\big),$
$\implies$ $\psi_{\theta}^{*}(m):=\sup_{\lambda\in\mathbb{R}}\{\lambda m-\psi_{\theta}(\lambda)\},$
where $\psi_{\theta}^{*}(m)$ is the Fenchel-Legendre transform of log-MGF. A symmetric argument with $\lambda<0$ yields the lower-tail bound
$\mathbb{P}_{\theta}(\bar{T}_{n}\leq m)\leq\exp\!\big(-n\psi_{\theta}^{*}(m)\big).$
We next link the Fenchel-Legendre transform to KL-Divergence using exponential tilting. First, adopting change of variables $\eta=\theta+\lambda$. Then
$\psi_{\theta}^{*}(m)$ $=\sup_{\eta}\{\langle\eta-\theta,m\rangle-(A(\eta)-A(\theta))\}$
$=A(\theta)+A^{*}(m)-\langle\theta,m\rangle,$
where $A^{*}(m)=\sup_{\eta}\{\langle\eta,m\rangle-A(\eta)\}$ is the convex conjugate of $A$. When $m$ is attainable, the supremum is achieved at $\eta=\theta_{m}$, and therefore
$\psi_{\theta}^{*}(m)=\langle\theta_{m}-\theta,m\rangle-(A(\theta_{m})-A(\theta)).$
But for exponential-family densities one has the following direct algebraic identity for the KL:
$D\big(p_{\theta_{1}}\|p_{\theta_{2}}\big)$ $=\mathbb{E}_{\theta_{1}}\Big[\log\frac{p_{\theta_{1}}(X)}{p_{\theta_{2}}(X)}\Big]$
$=\mathbb{E}_{\theta_{1}}\big[(\theta_{1}-\theta_{2})T(X)-(A(\theta_{1})-A(\theta_{2}))\big]$
$=(\theta_{1}-\theta_{2})\,\mathbb{E}_{\theta_{1}}[T(X)]-\big(A(\theta_{1})-A(\theta_{2})\big).$
Taking $\theta_{1}=\theta_{m}$ and $\theta_{2}=\theta$ (so $\mathbb{E}_{\theta_{1}}[T]=m$) yields
$\psi_{\theta}^{*}(m)=D\big(p_{\theta_{m}}\|p_{\theta}\big).$
Combining this with the one-sided Chernoff-bound above yields the one-sided KL-form Chernoff bounds
$\mathbb{P}_{\theta}(\bar{T}_{n}\geq m)\leq e^{-nD\bigl(p_{\theta_{m}}\|p_{\theta}\bigr)},$
$\mathbb{P}_{\theta}(\bar{T}_{n}\leq m)\leq e^{-nD\bigl(p_{\theta_{m}}\|p_{\theta}\bigr)}.$
Fix $\varepsilon>0$. Because $A^{\prime\prime}>0$ the function $m\mapsto D(p_{\theta_{m}}\|p_{\theta})$ is continuous, strictly convex and has a unique minimum 0 at $m=A^{\prime}(\theta)$. Thus the sublevel set $\{m:D(p_{\theta_{m}}\|p_{\theta})<\varepsilon\}$ is an open interval $(m_{-},m_{+})$; equivalently
$\{m:D(p_{\theta_{m}}\|p_{\theta})\geq\varepsilon\}=(-\infty,m_{-}]\cup[m_{+},\infty).$
Hence
$\{D(p_{\theta_{\overline{T}_{n}}}\|p_{\theta})\geq\varepsilon\}\subseteq\{\overline{T}_{n}\leq m_{-}\}\cup\{\overline{T}_{n}\geq m_{+}\}.$
Applying the one-sided KL bounds at $m_{\pm}$ (each equals $\varepsilon$) and using the union bound gives
$\mathbb{P}\bigl(D(p_{\theta_{\overline{T}_{n}}}\|p_{\theta})\geq\varepsilon\bigr)\leq e^{-n\varepsilon}+e^{-n\varepsilon}=2e^{-n\varepsilon},$
which proves the lemma.
Lemma A.4.
(Chernoff-Hoeffding Inequality)
[…]
$\bar{X}_{n}=\frac{1}{n}\sum_{i=1}^{n}X_{i}.$
Then for any $\varepsilon>0$,
$\mathbb{P}\left(D\left(\bar{X}_{n}\,\|\,\tilde{p}\right)>\varepsilon\right)\leq 2e^{-n\varepsilon},$
where $D(p\,\|\,q)$ is the Kullback-Leibler divergence between Bernoulli distributions with parameters $p$ and $q$.
Proof: Via Lemma A.3.
Lemma A.5 (Multinomial Chernoff-Hoeffding Bound via Mardia et al. (2019)).
For all $d\leq(\frac{nC_{0}}{4})^{\frac{1}{3}}$ and $P\in\mathcal{M}_{k}$, the following holds with the universal constants $C_{0}=\frac{e^{3}}{2\pi}\approx\;3.1967,$ for any $\epsilon>0$,
$\Pr\bigl(D(\widehat{P}_{n,d}\,\|\,P)\geq\epsilon\bigr)\leq 2(d-1)e^{-\frac{n\epsilon}{d-1}}.$
Lemma A.6 (Wasserstein distance bound via L.A. and Bhat (2022)).
Let $X$ be a sub-Gaussian r.v. with parameter $\sigma$. Let $F$ denote the CDF of $X$. Then, for every $n\geq 1$ and $\varepsilon$ such that $\tfrac{512\sigma}{\sqrt{n}}<\varepsilon<\tfrac{512\sigma}{\sqrt{n}}+16\sigma\sqrt{e}$, we have
$\mathbb{P}\big(W_{1}(F_{n},F)>\varepsilon\big)\;\leq\;\exp\!\left(-\frac{n}{256\sigma^{2}e}\,\Big(\varepsilon-\tfrac{512\sigma}{\sqrt{n}}\Big)^{2}\right),$
where $e$ is Euler’s number.
Lemma A.7 (Order-statistic thresholding).
Let $x_{1},\dots,x_{m}\in\mathbb{R}$ and let $x_{(1)}\leq\cdots\leq x_{(m)}$ be their order statistics. Fix a threshold $\tau\in\mathbb{R}$ and an integer $N\in\{1,\dots,m\}$. If at least $N$ of the sample values are at least $\tau$, i.e. $\big|\{j:\,x_{j}\geq\tau\}\big|\geq N$, then
$x_{(m-N+1)}\ \geq\ \tau.$
(Equivalently, if at least $N$ of the $x_{j}$ are strictly larger than $\tau$, the same conclusion holds.)
Proof:
Suppose, for contradiction, that $x_{(m-N+1)}<\tau$. Then all of the first mN+1m-N+1 order statistics are strictly less than $\tau$, so there are at most m(mN+1)=N1m-(m-N+1)=N-1 indices $j$ with $x_{j}\geq\tau$, contradicting $\big|\{j:\,x_{j}\geq\tau\}\big|\geq N$. Hence $x_{(m-N+1)}\geq\tau$.
Appendix B Proof to Theorem 3.1 and 5.1
B.1 Proof for Theorem 3.1
We condition on $\mathcal{G}_{j}:=\sigma(\psi_{j},\hat{q}_{j},n_{j})$ and $\mathcal{G}:=\sigma(\{\mathcal{G}_{j}\}_{j=1}^{m})$. By Assumption 3.1, conditional on $\mathcal{G}$ the additional randomness used to construct $\hat{p}_{j}$ (hence $\hat{\Delta}_{j}$ and the indicators below) is independent across $j$. For brevity, we abuse notation and write $p(\psi_{j}),q(\psi_{j})$ as $p_{j},q_{j}$, and similarly $\hat{p},\hat{q}$. We also ignore the superscript $k$ to simplify the proof, and discussions on the dependence on $k$ can be found in Section 3. Split $\delta=\delta_{\mathrm{DKW}}+\delta_{\mathrm{al}}+\delta_{\mathrm{cnt}}$ with $\delta_{\mathrm{DKW}}=\delta_{\mathrm{al}}=\delta_{\mathrm{cnt}}=\delta/3$, and union bound the three events defined below.
We first show an idealized case where we have access to the oracle $p_{j}$, and hence can obtain the actual sim-to-real discrepancy $\Delta_{j}$. Under this setting, we can directly define $F(t):=\mathbb{P}_{\psi\sim\Psi}(\Delta(\psi)\leq t)$ and
$\bar{F}_{m}(t)$ $:=\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\{\Delta_{j}\leq t\},$
$\bar{V}_{m}(u)$ $:=\inf\{t:\bar{F}_{m}(t)\geq u\}.$
By the DKW inequality (Lemma A.1), the following event
$\mathcal{E}_{\mathrm{DKW}}:=\Big\{\sup_{t}|\bar{F}_{m}(t)-F(t)|\leq\varepsilon_{m}(\delta)\Big\},$
occur with probability at least $1-\delta_{\mathrm{DKW}}$, where $\varepsilon_{m}(\delta)=\sqrt{\frac{\log(6/\delta)}{2m}}$. On $\mathcal{E}_{\mathrm{DKW}}$, for every $u\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta(\psi)\leq\bar{V}_{m}(u)\ \big|\ \mathcal{D}\big)\geq u-\varepsilon_{m}(\delta).$ (B.1)
Now in our dataset, we do not have access to the above oracle $\Delta_{j}$. Therefore, we control for $\Delta_{j}$ using the pseudo-discrepancy $\hat{\Delta}_{j}$. Define $Y_{j}:=\mathbf{1}\{\hat{\Delta}_{j}\geq\Delta_{j}\}.$ By the coverage of $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$ and the definition of $\hat{\Delta}_{j}$,
$\mathbb{P}(Y_{j}=1\mid\mathcal{G}_{j})\geq\gamma_{j}.$
For $r=1,\dots,m$, let $\alpha_{r}:=r/m$ and define the oracle right-tail set
$S_{r}:=\{j\in[m]:\Delta_{j}\geq\bar{V}_{m}(1-\alpha_{r})\},\quad s_{r}:=|S_{r}|.$
By definition of empirical quantiles, $s_{r}\geq r$. We then introduce link from the oracle set to the pseudo-discrepancies by defining counts
$T_{r}:=\sum_{j\in S_{r}}Y_{j},\quad U_{r}:=\bigl|\{j:\hat{\Delta}_{j}\geq\bar{V}_{m}(1-\alpha_{r})\}\bigr|.$
By construction, $U_{r}\geq T_{r}$. Notice that $U_{r}$ will give us the amount of $\hat{\Delta}$ required to achieve the tail count of the empirical oracle quantile function, hence we will try to create a lower bound for $U_{r}$ via $T_{r}$.
Under Assumption 3.1, $\{n_{j}\}$ are detereministic, and hence independent of $\{\psi_{j}\}$ and therefore independent of $\{\Delta_{j}\}$; consequently, the random index set $S_{r}$ (which is measurable w.r.t. $\{\Delta_{j}\}_{j=1}^{m}$) is independent of the vector $(\gamma_{1},\dots,\gamma_{m})$. Conditional on the realized multiset $\{\gamma_{j}\}_{j=1}^{m}$, the subcollection $\{\gamma_{j}:j\in S_{r}\}$ is thus a simple random sample without replacement from $\{\gamma_{j}\}_{j=1}^{m}$. By Hoeffding–Serfling Bardenet and Maillard (2015), for each $r$ and each $\tau>0$,
$\mathbb{P}\Big(\frac{1}{s_{r}}\sum_{j\in S_{r}}\gamma_{j}\leq\bar{\gamma}_{m}-\tau\Big)\leq\exp(-2s_{r}\tau^{2}).$
Set
$\tau_{r}:=\sqrt{\frac{\log(m/\delta_{\mathrm{al}})}{2s_{r}}}\leq\sqrt{\frac{\log(3m/\delta)}{2r}}=:\varepsilon_{\mathrm{al}}(\alpha_{r},\delta),$
and union bound over $r=1,\dots,m$ to obtain an event $\mathcal{E}_{\mathrm{al}}$ with $\mathbb{P}(\mathcal{E}_{\mathrm{al}})\geq 1-\delta_{\mathrm{al}}$ on which, for all $r$,
$\sum_{j\in S_{r}}\gamma_{j}\ \geq\ s_{r}\big(\bar{\gamma}_{m}-\varepsilon_{\mathrm{al}}(\alpha_{r},\delta)\big).$ (B.2)
Conditional on $\mathcal{G}$, the indicators $\{Y_{j}\}_{j=1}^{m}$ are independent and bounded in $[0,1]$, so Lemma A.2 implies that for any $t>0$,
$\mathbb{P}\Big(T_{r}\leq\mathbb{E}[T_{r}\mid\mathcal{G}]-t\Big|\mathcal{G}\Big)\leq\exp\Big(-\frac{2t^{2}}{s_{r}}\Big).$
Since $\mathbb{E}[T_{r}\mid\mathcal{G}]=\sum_{j\in S_{r}}\mathbb{E}[Y_{j}\mid\mathcal{G}]\geq\sum_{j\in S_{r}}\gamma_{j}$, choose
$c_{m}:=\sqrt{\frac{1}{2}\log\frac{m}{\delta_{\mathrm{cnt}}}}\leq\sqrt{\frac{1}{2}\log\frac{3m}{\delta}},\quad t_{r}:=c_{m}\sqrt{s_{r}}.$
Then $\mathbb{P}(T_{r}\geq\sum_{j\in S_{r}}\gamma_{j}-c_{m}\sqrt{s_{r}}\mid\mathcal{G})\geq 1-\delta_{\mathrm{cnt}}/m$. A union bound over $r=1,\dots,m$ yields an event $\mathcal{E}_{\mathrm{cnt}}$ with $\mathbb{P}(\mathcal{E}_{\mathrm{cnt}})\geq 1-\delta_{\mathrm{cnt}}$ on which, for all $r$,
$T_{r}\geq\sum_{j\in S_{r}}\gamma_{j}-c_{m}\sqrt{s_{r}}.$ (B.3)
Combining (B.2) and (B.3) gives, for all $r$,
$T_{r}\geq s_{r}\big(\bar{\gamma}_{m}-\varepsilon_{\mathrm{al}}(\alpha_{r},\delta)\big)-c_{m}\sqrt{s_{r}}$
on the event $\mathcal{E}_{\mathrm{al}}\cap\mathcal{E}_{\mathrm{cnt}}$. Using $s_{r}\geq r=m\alpha_{r}$ and $c_{m}/\sqrt{m}=b_{m}(\delta)$ yields
$U_{r}\geq T_{r}\geq m\Big(\big(\bar{\gamma}_{m}-\varepsilon_{\mathrm{al}}(\alpha_{r},\delta)\big)\alpha_{r}-b_{m}(\delta)\sqrt{\alpha_{r}}\Big).$
Equivalently, at least $m\alpha_{\mathrm{eff}}(\alpha_{r})$ of the $\hat{\Delta}_{j}$’s exceed $\bar{V}_{m}(1-\alpha_{r})$, hence
$\hat{V}_{m}\big(1-\alpha_{\mathrm{eff}}(\alpha_{r})\big)\geq\bar{V}_{m}(1-\alpha_{r}),\quad r=1,\dots,m.$ (B.4)
Finally, we work on $\mathcal{E}_{\mathrm{DKW}}\cap\mathcal{E}_{\mathrm{al}}\cap\mathcal{E}_{\mathrm{cnt}}$, which has probability at least $1-\delta$ by the union bound. Fix any $\alpha\in(0,1)$ and let $\alpha_{+}:=\lceil m\alpha\rceil/m$. By discretization, $\alpha_{\mathrm{eff}}(\alpha)=\alpha_{\mathrm{eff}}(\alpha_{+})$. Applying (B.4) at $r=\lceil m\alpha\rceil$ yields
$\hat{V}_{m}\big(1-\alpha_{\mathrm{eff}}(\alpha)\big)=\hat{V}_{m}\big(1-\alpha_{\mathrm{eff}}(\alpha_{+})\big)\geq\bar{V}_{m}(1-\alpha_{+}).$
Applying (B.1) with $u=1-\alpha_{+}$ gives
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta(\psi)\leq\bar{V}_{m}(1-\alpha_{+})\big|\mathcal{D}\big)$ $\geq 1-\alpha_{+}-\varepsilon_{m}(\delta)$
$\geq 1-\alpha-\varepsilon_{m}(\delta)-\frac{1}{m}.$
Monotonicity in the threshold implies (3.6), hence shown. ∎
Corollary B.1.
Adopt the setting of Theorem (3.1) and set $\gamma_{j}=\gamma$ for all $j\in[m]$. Then, for any $\alpha\in(0,1)$, with probability at least $1-\eta$ over $\mathcal{D}$, we have
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}^{(k)}\leq\hat{V}_{m}\!\left(1-\gamma\alpha\right)\Big|\mathcal{D}\Big)\geq 1-\alpha-\frac{\varepsilon_{\gamma}(\alpha,m,\eta)}{\sqrt{m}},$ (B.5)
where the remainder $\varepsilon_{\gamma}$ is $O\big(\sqrt{(\log m)/m}\big)$ as $m\to\infty$.
B.2 Proof for Corollary B.1:
One method is to directly adopt Theorem 3.1’s proof and readjust the remainder term. Here we adopt a slightly alternative method. We first consider a fixed level of confidence $\bar{\alpha}$, then extend to $\bar{\alpha}$ holding uniformly across $(0,1)$. Throughout the proof, $\alpha\in(0,1)$ denotes a generic quantile index (a function argument) and is distinct from the target coverage level $\bar{\alpha}$.
As a setup, we work conditionally on $\mathcal{G}_{j}:=\sigma(\psi_{j},\hat{q}_{j},n_{j}),\mathcal{G}:=\sigma\big(\{\mathcal{G}_{j}\}_{j=1}^{m}\big).$ Here $\mathcal{G}_{j}$ is the $\sigma$-field collecting all “human-side” information for scenario $j$. Thus $\mathcal{G}:=\sigma(\{\mathcal{G}_{j}\}_{j=1}^{m})$ represents the joint information from all scenarios that we treat as fixed when analyzing the remaining randomness coming from the estimation noise in $\hat{p}_{j}$. For brevity, we abuse notation and write $p(\psi_{j}),q(\psi_{j})$ as $p_{j},q_{j}$, and similarly $\hat{p},\hat{q}$. We also ignore the superscript $k$ to simplify the proof, and discussions on the dependence on $k$ can be found in Section 3. In addition, for any quantities $\{\Delta_{j}\}_{j=1}^{m}$, we denote the sorted version as $\{\Delta_{(i)}\}_{i=1}^{m}$, ie. $\Delta_{(1)}\leq\cdots\leq\Delta_{(m)}.$ For any sequence $\{\Delta_{j}\}_{j=1}^{m}$, let $\Delta_{(1)}\leq\cdots\leq\Delta_{(m)}$ denote its order statistics.
We seek the quantile function of $\Delta_{j}=L(p_{j},\hat{q}_{j})$, but only observe the estimators $(\hat{p}_{j},\hat{q}_{j})$. Note that the sequence $\{\Delta_{j}\}_{j=1}^{m}$ is i.i.d., since the simulator budget $k$ is fixed and each scenarios $\{\psi_{j}\}$ are i.i.d. as assumed in Section 2.
Let $\bar{V}_{m}(\alpha)=\inf\{t:\bar{F}_{m}(t)\geq\alpha\}=\Delta_{(\lceil m\alpha\rceil)}$ be the empirical $\alpha$ quantile of $\{\Delta_{j}\}_{j=1}^{m}$, where $\bar{F}_{m}(x)=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}(\Delta_{j}\leq x)$. By Lemma A.1, with probability $1-\delta$:
$\mathbb{P}_{\psi\sim\Psi}(L(p(\psi),\hat{q}(\psi))$ $\leq\bar{V}_{m}(1-\frac{\bar{\alpha}}{2})|\mathcal{D})$
$\geq 1-\frac{\bar{\alpha}}{2}-\sqrt{\frac{\log(2/\delta)}{2m}}$ (B.6)
and we would have our desired bound.
However, for the above argument to hold, we would need to know $p_{j}$, instead we only have the estimated $\hat{p}_{j}$. Therefore, we instead construct a pseudo-gaps using confidence sets with coverage level $\gamma$, and define the two gap terms as
$\begin{cases}{\Delta}_{j}=L(p_{j},\hat{q}_{j})&\text{, i.i.d., yet unobservable}\\ \hat{\Delta}_{j}:=\sup_{u\in C_{j}(\hat{p}_{j},\gamma)}L(u,\hat{q}_{j})&\text{, Not i.i.d., yet observable,}\end{cases}$
where $\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma\right)\subset\Theta$ are data-driven confidence sets satisfying $\mathbb{P}(p_{j}\in\mathcal{C}_{j}\!\left(\hat{p}_{j},\gamma\right)\big|\psi_{j},n_{j})\geq\gamma.$
By Assumption 3.2 and the compactness of the confidence set $C_{j}(\hat{p}_{j},\gamma)\subset\Theta$, Berge’s maximum theorem guarantees that the supremum in the definition of $\hat{\Delta}_{j}$ is attained, hence $\hat{\Delta}_{j}$ is well defined. Therefore, by the coverage property of $C_{j}$, we have $\mathbb{P}(\hat{\Delta}_{j}\geq{\Delta}_{j}|\mathcal{G}_{j}\cup\sigma(\gamma))\geq\gamma,$ and we get $\forall j$
$\mathbb{P}(\hat{\Delta}_{j}\geq{\Delta}_{j}|\mathcal{G}_{j})\geq\gamma.$ (B.7)
Furthermore, by the tower property and $\gamma\perp\!\!\!\perp\mathcal{G}_{j}$,
$\mathbb{P}\big(\hat{\Delta}_{j}\geq\Delta_{j}\big)=\mathbb{\mathbb{E}}\big[\mathbb{P}(\hat{\Delta}_{j}\geq\Delta_{j}\big|\mathcal{G}_{j})\big]\geq\gamma.$
Moreover, by the above discussion, conditional on $\mathcal{G}$, the indicators $Y_{j}=\mathbf{1}\{\hat{\Delta}_{j}\geq\Delta_{j}\}$ are independent across $j$. We will use $\{\hat{\Delta}_{j}\}_{j=1}^{m}$ to create an upper bound of $\bar{V}_{m}(1-\alpha)$, which along side (B.2) will give us our desired envelope. Intuitively, we want to find a larger quantile of $\hat{\Delta}_{j}$ and with (B.7), we can claim an upper bound of $\bar{V}_{m}(1-\alpha)$ with high probability.
First, define $S_{\alpha}:=\{j\in[m]:\Delta_{j}\geq\bar{V}_{m}(1-\alpha)\}$ and $s:=|S_{\alpha}|.$ By definition of the empirical $(1-\alpha)$–quantile $\bar{V}_{m}(1-\alpha)$, at most a fraction $(1-\alpha)$ of the sample can lie strictly below it. More precisely,
$\bigl|\{j\in[m]:\Delta_{j}<\bar{V}_{m}(1-\alpha)\}\bigr|\leq\bigl\lfloor m(1-\alpha)\bigr\rfloor.$
Therefore the number of indices with $\Delta_{j}\geq\bar{V}_{m}(1-\alpha)$ is at least
$s\geq m-\bigl\lfloor m(1-\alpha)\bigr\rfloor=\bigl\lceil m\alpha\bigr\rceil\geq\bigl\lfloor m\alpha\bigr\rfloor.$
We next define $Y_{j}=\mathbb{I}(\hat{\Delta}_{j}\geq\Delta_{j})$. By (B.7): $\mathbb{P}(Y_{j}=1|\mathcal{\mathcal{G}})\geq\gamma$. Next, define $T_{\alpha}=\{j\in S_{\alpha}:\hat{\Delta}_{j}\geq\Delta_{j}\}=\{j\in S_{\alpha}:Y_{j}=1\}=\{j:\hat{\Delta}_{j}\geq\Delta_{j}\geq\bar{V}_{m}(1-\alpha)\}$. First, we calculate $\mathbb{E}[T_{\alpha}]$:
$\mathbb{E}[T_{\alpha}|\mathcal{G}]=\mathbb{E}\Big[\sum_{j\in S_{\alpha}}Y_{j}|\mathcal{G}\Big]=\sum_{j\in S_{\alpha}}\mathbb{P}(Y_{j}|\mathcal{G})\geq\gamma s.$
Lemma A.2 implies that for any $\delta\in[0,\gamma s]$, we have:
$\mathbb{P}\Big(|T_{\alpha}|\leq\gamma s-t|\mathcal{G}\Big)$ $\leq\mathbb{P}\Big(|T_{\alpha}|\leq\mathbb{E}[T_{\alpha}]-t|\mathcal{G}\Big)$
$\leq\exp(-\frac{2t^{2}}{s}),$
where we applied $\mathbb{E}[T_{\alpha}|\mathcal{G}]\geq\gamma s$ in the first inequality. By setting $t=c\sqrt{s}$:
$\mathbb{P}\Big(|T_{\alpha}|\leq\gamma s-t|\mathcal{G}\Big)\leq\exp(-2c^{2})$ (B.8)
With this bound, we can link the actual set of indices we have interest, ie. $U_{\alpha}=\{j:\hat{\Delta}_{j}\geq\bar{V}_{m}(1-\alpha)\}$ to $T_{\alpha}$. By construction, $T_{\alpha}\subseteq U_{\alpha}$, hence by (B.8), with probability greather than $1-\exp(-2c^{2})$:
$|U_{\alpha}|\geq|T_{\alpha}|\geq\gamma s-c\sqrt{s},$
which implies at least $\gamma s-c\sqrt{s}$ of the $\hat{\Delta}_{j}$’s are larger than $\bar{V}_{m}(1-\alpha)$ with high probability.
We now analyze what coverage guarantee can we get for the inner probability via order statistics for any $\alpha$. Set $N:=\big\lfloor\gamma s-c\sqrt{s}\big\rfloor$ and define $\hat{V}_{m}(\alpha):=\inf\{t:\hat{F}_{m}(t)\geq\alpha\}=\Delta_{(\lceil m\alpha\rceil)}$, where $\hat{F}(x)=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}(\hat{\Delta}_{j}\leq x)$. If at least $N$ sample values exceed $\bar{V}_{m}(1-\alpha)$, then by order-statistics calculus (Lemma A.7)
$\hat{V}_{m}(1-\alpha)=\hat{\Delta}_{(m-\lfloor m\alpha\rfloor)}\geq\bar{V}_{m}(1-\alpha_{\mathrm{eff}}),$
whenever $\alpha_{\mathrm{eff}}$ is chosen so that $N\geq\lfloor m\alpha\rfloor+1$ holds when $s\geq\lfloor m\alpha_{\mathrm{eff}}\rfloor+1$.
A sufficient condition for the above to satisfy is
$\gamma m\alpha_{\mathrm{eff}}-c\sqrt{m\alpha_{\mathrm{eff}}}-1\geq m\alpha.$
Define $\alpha_{\mathrm{eff}}(\alpha,c,m):=\inf\{x\in(0,1):\gamma x-c\sqrt{\tfrac{x}{m}}-\tfrac{1}{m}\geq\alpha\}$. Writing $y=\sqrt{x}$, this is equivalent to $\gamma y^{2}-\tfrac{c}{\sqrt{m}}y-\big(\tfrac{1}{m}+\alpha\big)\geq 0$, so the minimal admissible $y$ is
$y^{*}$ $=\frac{c/\sqrt{m}+\sqrt{c^{2}/m+4\gamma\alpha+4\gamma/m}}{2\gamma},$
$\alpha_{\mathrm{eff}}(\alpha,c,m)$ $=(y^{*})^{2}.$
Thus, we define
$\mathcal{E}_{\alpha}$ $:=\Big\{\ \hat{V}_{m}(1-\alpha)\geq\bar{V}_{m}\big(1-\alpha_{\mathrm{eff}}(\alpha,c,m)\big)\Big\},$ (B.9)
$\text{then}\qquad\mathbb{P}(\mathcal{E}_{\alpha})\geq 1-e^{-2c^{2}}.$
Next, we apply (B.2) at level $1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)$:
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}\leq\bar{V}_{m}(1-\alpha_{\mathrm{eff}})\Big|\mathcal{D}\Big)$
$\geq 1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)-\varepsilon_{m}(\delta).$
On $\mathcal{E}_{\bar{\alpha}}$ in (B.9), $\bar{V}_{m}(1-\alpha_{\mathrm{eff}})\leq\hat{V}_{m}(1-\bar{\alpha})$. Hence,
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta_{\psi}\leq\hat{V}_{m}(1-\bar{\alpha})\big|\mathcal{D}\big)\geq 1-\alpha_{\mathrm{eff}}(\bar{\alpha},c,m)-\varepsilon_{m}(\delta),$ (B.10)
with outer probability at least $1-\delta-e^{-2c^{2}}$.
The exact algebraic form is
$\alpha_{\mathrm{eff}}(\bar{\alpha},$ $c,m)=\frac{\big(\tfrac{c}{\sqrt{m}}+\sqrt{\tfrac{c^{2}}{m}+4\gamma\bar{\alpha}+\tfrac{4\gamma}{m}}\big)^{2}}{4\gamma^{2}}=$
$\frac{\bar{\alpha}}{\gamma}+\frac{c}{\gamma\sqrt{m}}\sqrt{4\gamma\bar{\alpha}+\tfrac{c^{2}+4\gamma}{m}}+\frac{c^{2}}{2\gamma^{2}m}+\frac{1}{\gamma m}.$
Therefore, as $m\to\infty$, $\alpha_{\mathrm{eff}}(\alpha,c,m)=\tfrac{\bar{\alpha}}{\gamma}+c\sqrt{4\bar{\alpha}/\gamma m}+O(m^{-1})$.
We have shown that for any target level $\alpha$ and choice of $c>0$, the preceding argument yields a high-probability concentration bound based on $\{\hat{\Delta}_{j}\}_{j=1}^{m}$. We now extend the guarantee to hold uniformly over all $\alpha$. Fix $c>0$ and $\delta\in(0,1)$. For the grid $\alpha_{r}:=r/m$ ($r=1,\dots,m$), let
$\mathcal{E}_{\mathrm{DKW}}$ $:=\Big\{\sup_{x}\big|\hat{F}_{m}(x)-F^{*}(x)\big|\leq\varepsilon_{m}(\delta)\Big\}$
[…]
By DKW, $\mathbb{P}(\mathcal{E}_{\mathrm{DKW}})\geq 1-\delta$, and by the fixed-level argument, $\mathbb{P}(\mathcal{E}_{r})\geq 1-e^{-2c^{2}}$ for each $r$. Hence, by a union bound,
$\mathbb{P}\Big(\mathcal{E}_{\mathrm{DKW}}\cap\bigcap_{r=1}^{m}\mathcal{E}_{r}\Big)\geq 1-\delta-me^{-2c^{2}}.$
For any $\alpha\in(0,1)$ let $r=\lceil m\alpha\rceil$ and denote $\alpha_{+}:=\alpha_{r}=r/m\in[\alpha,\alpha+1/m]$. Since the empirical quantile is piecewise constant on the $m$-grid,
$\widehat{V}_{m}(1-\alpha)=\widehat{V}_{m}(1-\alpha_{+}).$
Applying (B.10) at level $\alpha_{+}$ yields
$\mathbb{P}_{\psi\sim\Psi}$ $\Big(\Delta(\psi)\leq\widehat{V}_{m}(1-\alpha)\Big|\mathcal{D}\Big)$
$\geq 1$ $-\frac{\alpha_{+}}{\gamma}-\frac{c}{\gamma\sqrt{m}}\sqrt{4\gamma\alpha_{+}+\tfrac{c^{2}+4\gamma}{m}}$
$-\frac{c^{2}}{2\gamma^{2}m}-\frac{1}{\gamma m}-\varepsilon_{m}(\delta).$
Since $\alpha_{+}\in[\alpha,\alpha+1/m]$ and the right-hand side is nonincreasing in $\alpha$, the same bound holds with $\alpha$ replaced by $\alpha_{+}$, and (optionally) one may absorb the rounding slack $\alpha_{+}-\alpha\leq 1/m$ into the $O(m^{-1})$ term by a crude inequality
$\frac{\alpha_{+}}{\gamma}\leq\frac{\alpha}{\gamma}+\frac{1}{\gamma m},\quad\sqrt{4\gamma\alpha_{+}+\tfrac{c^{2}+4\gamma}{m}}\leq\sqrt{4\gamma\alpha+\tfrac{c^{2}+8\gamma}{m}}.$
Therefore, with probability at least $1-\delta-me^{-2c^{2}}$ over $\mathcal{D}$, the guarantee (B.5) (at $\alpha$ replaced by $\alpha_{+}$) holds uniformly for all $\alpha\in(0,1)$. The form in the main theorem is a simplification with respect to $c$ and $\delta$ while taking $\gamma=\tfrac{1}{2}$.
B.3 Corollary for simulator budget:
Corollary B.2.
Suppose Assumptions 3.1 and 3.2 hold. For any simulation sample size $k_{j}\in\mathbb{N}$, define the per-scenario true sim-to-real discrepancy and its pseudo-discrepancy by
$\Delta_{j}^{\star,(k)}:=L(p_{j},q_{j}),\qquad\hat{\Delta}_{j}^{\star,(k)}:=\sup_{\begin{subarray}{c}u\in C_{j}^{p}(\hat{p}_{j})\\ v\in C_{j}^{q}(\hat{q}_{j})\end{subarray}}L(u,v),$
where $C_{j}^{p}(\hat{p}_{j})\subset\Theta_{p}$, $C_{j}^{q}(\hat{q}_{j})\subset\Theta_{q}$ are data-driven compact confidence sets such that
$\mathbb{P}\big(p_{j}\in C_{j}^{p}(\hat{p}_{j}),\ q_{j}\in C_{j}^{q}(\hat{q}_{j})\big|\psi_{j},n_{j},k_{j}\big)\geq\tfrac{1}{2}.$
Let $\hat{V}_{m}^{\star}(\alpha)$ denote the empirical $\alpha$-quantile of $\{\hat{\Delta}_{j}^{\star,(k)}\}_{j=1}^{m}$.
Then, for any $\alpha\in(0,1)$ and any $\eta\in(0,1)$, with probability at least $1-\eta$ over $\mathcal{D}$, we have
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta_{\psi}^{\star,(k)}\leq\hat{V}_{m}^{\star}\big(1-\tfrac{\alpha}{2}\big)\Big|\mathcal{D}\Big)\geq 1-\alpha-\frac{\varepsilon(\alpha,m,\eta)}{\sqrt{m}},$ (B.11)
where the remainder $\varepsilon(\alpha,m,\eta)$ is the same as in Theorem B.1, in particular $\varepsilon(\alpha,m,\eta)=O\big(\sqrt{\log m}\big)$ and $\varepsilon(\alpha,m,\eta)/\sqrt{m}=O\big(\sqrt{(\log m)/m}\big)$.
Proof.
The argument is identical to the proof of Theorem B.1, except that we now construct separate confidence sets for $p_{j}$ and $q_{j}$ with marginal coverages $\gamma_{p}$ and $\gamma_{q}$ chosen so that $\gamma_{p}\gamma_{q}=\tfrac{1}{2}$. For example, we can set $\gamma_{p}=\gamma_{q}=\tfrac{1}{\sqrt{2}}$. This ensures that the joint event $\{p_{j}\in C_{j}^{p}(\hat{p}_{j}),\,q_{j}\in C_{j}^{q}(\hat{q}_{j})\}$ plays exactly the same role as the univariate coverage event in Theorem 3.1, i.e., $\mathbb{P}(\hat{\Delta}_{j}^{\star,(k)}\geq\Delta_{j}^{\star,(k)}\mid\mathcal{G}_{j})\geq\tfrac{1}{2}$. All subsequent steps then carry over verbatim, yielding the stated bound. ∎
Theorem B.1.
Suppose the setup of Theorem 3.1 and Assumptions 3.13.2 hold. Let $\{\gamma_{L,j},\gamma_{U,j}\}_{j=1}^{m}\subset(0,1]$ be coverage levels and define the lower and upper pseudo-discrepancies
$\hat{\Delta}_{j}^{-}:=\inf_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{L,j})}L(u,\hat{q}_{j}),\ \hat{\Delta}_{j}^{+}:=\sup_{u\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{U,j})}L(u,\hat{q}_{j}),$
where $\mathcal{C}_{j}(\hat{p}_{j},\gamma)\subset\Theta$ are data-driven compact confidence sets satisfying the conditional coverage guarantees
$\mathbb{P}\big(p_{j}\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{L,j})\mid\psi_{j},n_{j}\big)$ $\geq\gamma_{L,j}$
$\mathbb{P}\big(p_{j}\in\mathcal{C}_{j}(\hat{p}_{j},\gamma_{U,j})\mid\psi_{j},n_{j}\big)$ $\geq\gamma_{U,j}.$
Let $\hat{V}_{m}^{-}(\cdot)$ denote the empirical quantile function of $\{\hat{\Delta}_{j}^{-}\}_{j=1}^{m}$, and $\hat{V}_{m}^{+}(\cdot)$ the empirical quantile function of $\{\hat{\Delta}_{j}^{+}\}_{j=1}^{m}$. Define the empirical mean coverage levels $\bar{\gamma}_{L,m}:=\frac{1}{m}\sum_{j=1}^{m}\gamma_{L,j},\bar{\gamma}_{U,m}:=\frac{1}{m}\sum_{j=1}^{m}\gamma_{U,j}.$ For $\alpha\in(0,1)$ and $\delta\in(0,1)$ define the discretized level $\tilde{\alpha}(\alpha):=\frac{\lceil m\alpha\rceil}{m}.$ Define the effective indices
$\alpha_{\mathrm{eff}}^{L}(\alpha)$ $:=\Big[\big(\bar{\gamma}_{L,m}-\varepsilon_{\mathrm{al}}(\alpha,\delta)\big)\tilde{\alpha}(\alpha)-b_{m}(\delta)\sqrt{\tilde{\alpha}(\alpha)}\Big]_{+},$
$\alpha_{\mathrm{eff}}^{U}(\alpha)$ $:=\Big[\big(\bar{\gamma}_{U,m}-\varepsilon_{\mathrm{al}}(\alpha,\delta)\big)\tilde{\alpha}(\alpha)-b_{m}(\delta)\sqrt{\tilde{\alpha}(\alpha)}\Big]_{+},$
where the remainder terms are defined as $\varepsilon_{\mathrm{al}}(\alpha,\delta):=\sqrt{\frac{\log(3m/\delta)}{2m\tilde{\alpha}(\alpha)}},$ and $b_{m}(\delta):=\sqrt{\frac{1}{2m}\log\frac{3m}{\delta}}.$ Then, with probability at least $1-\delta$ over $\mathcal{D}$, simultaneously for all $\alpha\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}\big(\alpha_{\mathrm{eff}}^{L}(\alpha)\big)\Big|\mathcal{D}\Big)$ $\geq 1-\alpha-\varepsilon_{m}(\delta),$
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\leq\hat{V}_{m}^{+}\big(1-\alpha_{\mathrm{eff}}^{U}(\alpha)\big)\Big|\mathcal{D}\Big)$ $\geq 1-\alpha-\varepsilon_{m}(\delta),$
where $\varepsilon_{m}(\delta):=\sqrt{\frac{\log(6/\delta)}{2m}}+\frac{1}{m}$. Let $V:[0,1]\to\mathbb{R}$ denote the quantile function of $\Delta(\psi)$. In particular, for any fixed $\tau\in(0,1)$, as $m\to\infty$, the band admits the heuristic form
$\hat{V}_{m}^{-}\big(\bar{\gamma}_{L,m}\tau\big)\lesssim V(\tau)\lesssim\hat{V}_{m}^{+}\big(\bar{\gamma}_{U,m}\tau+(1-\bar{\gamma}_{U,m})\big)$ (B.12)
B.4 Proof for Theorem B.1:
We follow the setup of Theorem 3.1 and only provide the lower bound proof, since upper bound is exactly Theorem 3.1. Recall that we work conditionally on $\mathcal{G}_{j}:=\sigma(\psi_{j},\hat{q}_{j},n_{j})$ and $\mathcal{G}:=\sigma\big(\{\mathcal{G}_{j}\}_{j=1}^{m}\big)$. For brevity, write $p_{j}:=p(\psi_{j})$, $q_{j}:=q(\psi_{j})$, and similarly $\hat{p}_{j},\hat{q}_{j}$. The true per-scenario discrepancy is $\Delta_{j}:=L(p_{j},\hat{q}_{j})$, $j=1,\dots,m$, and by the i.i.d. scenario assumption, $\{\Delta_{j}\}_{j=1}^{m}$ are i.i.d. draws from the distribution of $\Delta(\psi)$ under $\psi\sim\Psi$.
Define the lower pseudo-discrepancies
$\Delta_{j}^{-}:=\inf_{u\in\mathcal{C}_{j}(\hat{p}_{j})}L(u,\hat{q}_{j}),\qquad j=1,\dots,m,$
where $\mathcal{C}_{j}(\hat{p}_{j})\subset\Theta$ satisfy the lower coverage guarantee $\mathbb{P}\big(p_{j}\in\mathcal{C}_{j}(\hat{p}_{j})\mid\psi_{j},n_{j}\big)\geq\gamma_{L}.$ By Assumption 3.2 and compactness of $\mathcal{C}_{j}(\hat{p}_{j})$, Berge’s maximum theorem implies that the infimum is attained on $\mathcal{C}_{j}(\hat{p}_{j})$, hence $\Delta_{j}^{-}$ is well defined.
Also by definition, on the event $\{p_{j}\in\mathcal{C}_{j}(\hat{p}_{j})\}$ we have $\Delta_{j}^{-}\leq\Delta_{j}$. Thus, for each $j$,
$\mathbb{P}\big(\Delta_{j}^{-}\leq\Delta_{j}\big|\mathcal{G}_{j}\big)\geq\mathbb{P}\big(p_{j}\in\mathcal{C}_{j}(\hat{p}_{j})\big|\mathcal{G}_{j}\big)\geq\gamma_{L}.$
Define the indicators $Y_{j}^{-}:=\mathbf{1}\{\Delta_{j}^{-}\leq\Delta_{j}\}$, $j=1,\dots,m$. Conditionally on $\mathcal{G}$, the variables $\{Y_{j}^{-}\}_{j=1}^{m}$ are independent and satisfy
$\mathbb{P}(Y_{j}^{-}=1\mid\mathcal{G})\geq\gamma_{L},\qquad\forall j.$ (B.13)
We will fix $\alpha\in(0,1)$ for the moment, then generalize toward $\forall\alpha$. We again denote $\bar{V}_{m}(\alpha)$ as the empirical $\alpha$-quantile of the (unobservable) $\{\Delta_{j}\}_{j=1}^{m}$, i.e., $\bar{V}_{m}(\alpha):=\Delta_{(\lceil m\alpha\rceil)}$, where $\Delta_{(1)}\leq\cdots\leq\Delta_{(m)}$ are the order statistics, and let $\bar{F}_{m}(t):=\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\{\Delta_{j}\leq t\}$ be their empirical CDF. Define the index set of “lower-tail” scenarios
$S_{\alpha}^{-}:=\{j\in[m]:\Delta_{j}\leq\bar{V}_{m}(\alpha)\},\quad s^{-}:=|S_{\alpha}^{-}|.$
By definition of $\bar{V}_{m}(\alpha)$ we have $s^{-}\geq\lceil m\alpha\rceil$. Among these, consider those for which the lower pseudo-gap is valid, where we denote $T_{\alpha}^{-}:=\{j\in S_{\alpha}^{-}:Y_{j}^{-}=1\}=\{j\in S_{\alpha}^{-}:\Delta_{j}^{-}\leq\Delta_{j}\leq\bar{V}_{m}(\alpha)\}.$ By (B.13),
$\mathbb{E}\big[|T_{\alpha}^{-}|\big|\mathcal{G}\big]=\sum_{j\in S_{\alpha}^{-}}\mathbb{P}(Y_{j}^{-}=1\mid\mathcal{G})\geq\gamma_{L}s^{-}.$
Applying Lemma A.2, for any $t\in[0,\gamma_{L}s^{-}]$,
$\mathbb{P}\Big(|T_{\alpha}^{-}|\leq\gamma_{L}s^{-}-t\Big|\mathcal{G}\Big)\leq\exp\Big(-\frac{2t^{2}}{s^{-}}\Big).$
Setting $t=c\sqrt{s^{-}}$ for some $c>0$ yields
$\mathbb{P}\Big(|T_{\alpha}^{-}|\leq\gamma_{L}s^{-}-c\sqrt{s^{-}}\Big|\mathcal{G}\Big)\leq e^{-2c^{2}}.$ (B.14)
Define also $U_{\alpha}^{-}:=\{j\in[m]:\Delta_{j}^{-}\leq\bar{V}_{m}(\alpha)\}.$ Then $T_{\alpha}^{-}\subseteq U_{\alpha}^{-}$, so on the complement of the event in (B.14),
$|U_{\alpha}^{-}|\geq|T_{\alpha}^{-}|\geq\gamma_{L}s^{-}-c\sqrt{s^{-}}.$
We now analyze the guarantee we can get for the inner probability via order statistics. Set $N:=\big\lfloor\gamma_{L}s^{-}-c\sqrt{s^{-}}\big\rfloor$ on the complement of the event in (B.14), so that $|U_{\alpha}^{-}|\geq N$ holds. Let $\hat{F}_{m}^{-}(t)$ denote the empirical CDF of $\{\Delta_{j}^{-}\}_{j=1}^{m}$, $\hat{F}_{m}^{-}(t):=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}(\Delta_{j}^{-}\leq t),$ and define the empirical $\beta$-quantile $\hat{V}_{m}^{-}(\beta):=\inf\{t:\hat{F}_{m}^{-}(t)\geq\beta\}=\Delta^{-}_{(\lceil m\beta\rceil)}.$ If at least $N$ sample values satisfy $\Delta_{j}^{-}\leq\bar{V}_{m}(\alpha)$, then by Lemma A.7,
$\hat{V}_{m}^{-}(\beta)=\Delta^{-}_{(\lceil m\beta\rceil)}\leq\bar{V}_{m}(\alpha),$
whenever $\lceil m\beta\rceil\leq N$, i.e. whenever $\beta\leq N/m$.
A sufficient condition for $\beta\leq N/m$ is obtained by lower bounding $N/m$ using $s^{-}\geq\lceil m\alpha\rceil$. Indeed, on the same event,
$\frac{N}{m}\geq\frac{\gamma_{L}s^{-}-c\sqrt{s^{-}}-1}{m}\geq\gamma_{L}\alpha-c\sqrt{\frac{\alpha}{m}}-\frac{1}{m},$
where the last inequality uses $s^{-}\geq\lceil m\alpha\rceil\geq m\alpha$ and the floor/ceiling slack. Define
$\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m):=\gamma_{L}\alpha-c\sqrt{\frac{\alpha}{m}}-\frac{1}{m}.$ (B.15)
Then the preceding display implies $\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\leq N/m$, and hence
$\hat{V}_{m}^{-}\Big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\Big)\ \leq\ \bar{V}_{m}(\alpha)$ (B.16)
on the complement of the event in (B.14).
Thus, we define
$\mathcal{E}_{\alpha}^{-}$ $:=\Big\{\ \hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\ \leq\ \bar{V}_{m}(\alpha)\ \Big\},$ (B.17)
$\text{then}\qquad\mathbb{P}(\mathcal{E}_{\alpha}^{-})\ \geq\ 1-e^{-2c^{2}}.$
Next, we bridge $\bar{F}_{m}$ to the population CDF of $\Delta(\psi)$ under $\psi\sim\Psi$, which we denote by $F_{\Delta}$. By Lemma A.1, for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\mathcal{E}_{\mathrm{DKW}}:=\Big\{\sup_{t\in\mathbb{R}}\big|\bar{F}_{m}(t)-F_{\Delta}(t)\big|\leq\varepsilon_{m}(\delta)\Big\},$
where $\varepsilon_{m}(\delta):=\sqrt{\frac{\log(2/\delta)}{2m}}$.
Work on the intersection $\mathcal{E}_{\mathrm{DKW}}\cap\mathcal{E}_{\alpha}^{-}$. By (B.17) and the monotonicity of $F_{\Delta}$,
$F_{\Delta}\Big(\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\Big)\leq F_{\Delta}\big(\bar{V}_{m}(\alpha)\big).$
On $\mathcal{E}_{\mathrm{DKW}}$, we have
$F_{\Delta}\big(\bar{V}_{m}(\alpha)\big)\leq\bar{F}_{m}\big(\bar{V}_{m}(\alpha)\big)+\varepsilon_{m}(\delta)\leq\alpha+\varepsilon_{m}(\delta),$
where the last inequality uses $\bar{F}_{m}(\bar{V}_{m}(\alpha))\geq\alpha$ by definition of $\bar{V}_{m}(\alpha)$. Combining the last two displays yields
$F_{\Delta}\Big(\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\Big)\leq\alpha+\varepsilon_{m}(\delta).$
Equivalently,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\Big|\mathcal{D}\Big)\geq 1-\alpha-\varepsilon_{m}(\delta),$ (B.18)
with outer probability at least $1-\delta-e^{-2c^{2}}$.
We have shown that for any target level $\alpha$ and choice of $c>0$, the preceding argument yields a high-probability bound based on $\{\Delta_{j}^{-}\}_{j=1}^{m}$. We now extend the guarantee to hold uniformly over all $\alpha$. Fix $c>0$ and $\delta\in(0,1)$. For the grid $\alpha_{r}:=r/m$ ($r=1,\dots,m$), let
$\mathcal{E}_{\mathrm{DKW}}$ $:=\Big\{\sup_{t}\big|\bar{F}_{m}(t)-F_{\Delta}(t)\big|\leq\varepsilon_{m}(\delta)\Big\},$
[…]
By DKW, $\Pr(\mathcal{E}_{\mathrm{DKW}})\geq 1-\delta$, and by the fixed-level argument, $\Pr(\mathcal{E}_{r}^{-})\geq 1-e^{-2c^{2}}$ for each $r$. Hence, by a union bound,
$\mathbb{P}\Big(\mathcal{E}_{\mathrm{DKW}}\cap\bigcap_{r=1}^{m}\mathcal{E}_{r}^{-}\Big)\geq 1-\delta-me^{-2c^{2}}.$
For an arbitrary $\alpha\in(0,1)$, let $r=\lceil m\alpha\rceil$ and denote $\alpha_{+}:=\alpha_{r}=r/m\in[\alpha,\alpha+1/m]$. Applying (B.18) at $\alpha_{+}$ gives
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha_{+},c,m)\big)\Big|\mathcal{D}\Big)\geq 1-\alpha_{+}-\varepsilon_{m}(\delta).$
Since $\alpha_{+}\geq\alpha$, we have $\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha_{+},c,m)\geq\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)$, and because $\hat{V}_{m}^{-}(\cdot)$ is nondecreasing,
$\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\leq\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha_{+},c,m)\big).$
Therefore,
$\mathbb{P}_{\psi\sim\Psi}\Big(\Delta(\psi)\geq\hat{V}_{m}^{-}\big(\beta_{\mathrm{eff}}^{-,(\gamma_{L})}(\alpha,c,m)\big)\Big|\mathcal{D}\Big)\geq 1-\alpha_{+}-\varepsilon_{m}(\delta),$
and one may absorb the rounding slack $\alpha_{+}-\alpha\leq 1/m$ into the $O(m^{-1})$ term (as in Theorem 3.1). Hence, with probability at least $1-\delta-me^{-2c^{2}}$ over $\mathcal{D}$, the bound (B.18) holds uniformly for all $\alpha\in(0,1)$ (up to the $1/m$ discretization slack). This completes the proof.
Appendix C World Value Bench Methodology
C.1 Selection of Survey Questions
The World Values Survey contains 259 questions, grouped into categories such as social values, well-being, economic values, and security. We exclude questions that are difficult to interpret along an ordered sentiment scale. For example, in Figure 8, Question 223 is highly country- and time-specific and does not admit a natural ordering of sentiment, making it hard to align with other questions.
Figure 8: Text of Question 223.
After this screening, we retain 235 questions, each with more than 90,000 responding participants. Ideally, we would restrict the question pool to questions within a single category, but in practice we prioritize having a sufficiently large number of questions. Consequently, we keep all retained questions when conducting our experiments.
C.2 Example Question and Preprocess
Below we list three example questions from the dataset.
Figure 9: Example questions from the Political Interest category.
Figure 10: Example question from the Science and Technology category.
Figure 11: Example question from the Migration category.
As discussed in the main text, to obtain a unified scale across heterogeneous categorical response formats, we map all answers to numerical values in [1,1][-1,1]. Specifically, we query GPT-5 to determine the direction of this mapping according to its assessment of the “idealness” of each response in terms of public opinion. For example, in Figure 10, GPT-5 maps option 1 (“A lot worse off”) to -1 and option 10 (“A lot better off”) to 1, with intermediate choices linearly spaced between these endpoints. [^3] This conversion does not represent the authors’ values; it reflects GPT-5’s inferred public opinion, which may be influenced by RLHF or other factors unknown to the authors.
C.3 Synthetic Profile Generation
In the original dataset, some respondent IDs appear irregularly or are duplicated. We exclude these records and are left with 96,220 unique participants. The dataset contains demographic information for each participant, summarized in Table 2.
Table 2: Master list of demographic variables and categories.
Demographic Categories
Sex Male; Female
Country 64 countries/societies
Age Numeric (years)
Language Text
Migration status Native-born; First-generation immigrant; Second-generation; Other/missing
Marital status Married; Living together as married; Divorced; Separated; Widowed; Never married/Single
Children status Number of children (numeric)
Education level ISCED levels 0–8444ISCED 0 corresponds to early childhood or no education; ISCED 8 corresponds to doctoral or equivalent.
Employment status Employed full-time; Employed part-time; Self-employed; Unemployed; Student; Homemaker; Retired; Other
Household income decile 1 (lowest) ; 2; 3; 4; 5; 6; 7; 8; 9; 10 (highest)
Religious affiliation Christian; Muslim; Hindu; Buddhist; Jewish; No religion; Other religion
Self-identified social class Lower class; Working class; Lower middle class; Upper middle class; Upper class
Settlement Type Urban; Rural (with Population Size)
We select a subset of these demographic attributes to define our synthetic profiles, including Country, Sex, Age, and Settlement Type, among others. The LLMs are asked to produce a categorical response encoded as [[1]], [[2]], etc. For all models, we prepend the following instruction when querying the API: “You are simulating the behaviors of humans with certain specified characteristics to help with a survey study.” We illustrate the resulting query structure in the example prompt below.
Example prompt: Pretend that you reside in Mexico. You live in a capital city urban with a town size of 100,000-500,000 area. You are female, your age is between 35-44. and you are Married. You normally speak Spanish; Castilian at home. In terms of migration background, you were born in this country (not an immigrant). In terms of education, you attained Upper secondary education (ISCED 3). Your current employment status is: Housewife not otherwise employed. You work in Not applicable; Never had a job. You belong to Roman Catholic; Latin Church;. On a scale of 1 to 10, 1 meaning ’A lot worse off’ and 10 meaning ’A lot better off’, do you think science and technology have improved or worsened the world? Please respond with a single number from 1 to 10 in double square brackets, e.g., [[1]].
C.4 Adaptive coverage rate $\gamma_{j}$ choice
As mentioned in the main text, we focus on the case where we set the coverage rate to the generic functional form
$\gamma_{j}=1-n_{j}^{-\beta}.$
We consider a set of $\beta$ values in our sensitivity analysis. We conduct a numerical experiment over $\beta\in\{\tfrac{1}{5},\tfrac{1}{4},\tfrac{1}{3},\tfrac{1}{2}\}$ in Figure 12. The resulting fidelity profiles are largely stable across these $\gamma_{j}$ schedules, and we adopt $\beta=\tfrac{1}{3}$ in the main text for simplicity.
Figure 12: $\beta$ sensitivtiy analysis aunder $n=\{50,200\}$.
Appendix D Simulation System Examples
Manufacturing: Factory Production (discrete-event simulation; cycle time).
  • Outcome space: $\mathcal{X}=\mathbb{R}_{+}$ (cycle time or throughput).
  • Scenarios: $\psi$ = product mix + scheduling policy.
  • Profiles: $\mathcal{Z}$ = machine/operator states, shift team, lot sizes; $\mathcal{P}$ = plant variability.
  • Laws: $Q^{\mathrm{gt}}(\cdot\mid z,\psi)$ = empirical cycle-time distribution on the floor; $Q^{\mathrm{sim}}(\cdot\mid a(z),\psi,r)$ = DES output under mirrored inputs.
  • Parameters: $\Theta=\mathbb{R}$ (mean cycle time) or $\mathbb{R}^{2}$ (mean, variance).
  • Discrepancy: $L$ = difference of means, Gaussian KL.
  • Sampling: $n_{j}$ production runs logged; $k$ simulated replications per scenario.
Environment: Urban decarbonization (technology choice; multinomial).
  • Outcome space: $\mathcal{X}=\{1,\dots,K\}$ with mean $p(\psi)\in\Delta^{K-1}$ (for example, gas furnace, heat pump, variable refrigerant flow, other).
  • Scenarios: $\psi$ consists of city, season, rebate level, carbon price path, and policy bundle.
  • Profiles: $\mathcal{Z}$ contains household and building attributes such as income, occupants, roof area, and baseline electricity use, or exogenous drivers including weather and demand shocks.
  • Laws: $Q^{\mathrm{gt}}(\cdot\mid z,\psi)$ denotes the empirical technology-choice distribution, and $Q^{\mathrm{sim}}(\cdot\mid a(z),\psi,r)$ denotes the simulator output under mirrored inputs.
  • Parameters: $\Theta=\Delta^{K-1}$ for category probabilities, or a low-dimensional reparameterization such as multinomial logistic parameters.
  • Discrepancy: $L$ on the simplex, for example the total-variation distance $\tfrac{1}{2}\lVert p-q\rVert_{1}$, the multiclass Kullback–Leibler divergence $\sum_{c=1}^{K}p_{c}\log\!\bigl(p_{c}/q_{c}\bigr)$.
  • Sampling: $n_{j}$ human records per scenario and $k$ synthetic replications per scenario.
Appendix E Applications
E.1 EEDI Dataset
Our dataset is EEDI He-Yueya et al. (2024), built on the NeurIPS 2020 Education Challenge Wang et al. (2021), which consists of student responses to mathematics multiple-choice questions collected on the Eedi online education platform. The full corpus includes 573 distinct questions and 443,433 responses from 2,287 students, and each question has four options A-D that we binarize as “correct/incorrect” based on the student’s or simulator’s choice, consistent with Lemma A.4. We adopt the preprocessed version curated by Huang et al. (2025), which retains questions with at least 100 student responses and excludes items with graphs or diagrams, yielding 412 questions. EEDI also provides individual-level covariates such as gender, age, and socioeconomic status, which the authors of Huang et al. (2025) use to construct synthetic profiles. Under the same problem formulation, they compute $\{\hat{p}_{j},\hat{q}_{j}\}_{j=1}^{412}$ for seven LLMs: GPT-3.5-Turbo (gpt-3.5-turbo), GPT-4o (gpt-4o), and GPT-4o-mini (gpt-4o-mini); Claude 3.5 Haiku (claude-3-5-haiku-20241022); Llama 3.3 70B (Llama-3.3-70B-Instruct-Turbo); Mistral 7B (Mistral-7B-Instruct-v0.3); DeepSeek-V3 (DeepSeek-V3), and constructed a benchmark random simulator that selects uniformly among the available answer choices. A more detailed exploration into the EEDI dataset and the simulation procedure can be found in Huang et al. (2025).
We apply our methodology to produce a fidelity profile for each candidate LLM. We use absolute error as the loss, L(p,q)=|pq|L(p,q)=|p-q| and set the simulation budget $k=50$.
Figure 13: Quantile fidelity profiles $\hat{V}(\alpha)$ across LLMs (Discrepancy: Absolute loss, $k=50$.
Figure 13 compares models by how tightly their synthetic outcomes track the human distribution across items. We plot $\hat{V}_{\ell}(\alpha)$ against $\alpha$, where lower-flatter curves indicate uniformly small discrepancies, while elbows reveal rare but severe misses. DeepSeek-V3 lies lowest across most quantiles, indicating the most reliable alignment, with the random benchmark and GPT-4o close behind. Notably, several models do not outperform the random baseline, suggesting they may be ill-suited for agent-based simulation under this discrepancy function.
E.2 OpinionQA
Our dataset is OpinionQA Santurkar et al. (2023b), built from the Pew Research’s American Trends Panel, which consists of the US population’s responses to survey questions spanning topics such as racial equity, security, and technology. We adopt the preprocessed version curated by Huang et al. (2025), which includes 385 distinct questions and 1,476,868 responses from at least 32,864 people. Each question has 5 choices, corresponding to the order sentiments, which is a multinomial setting. We can construct confidence sets $\mathcal{C}_{j}$ for multinomial vectors by adopting Example 3.1 with $d=5$. OpinionQA also provides individual-level covariates such as gender, age, socioeconomic status, religious affiliation, and marital status, and more, which are used to construct synthetic profiles. Under the same problem formulation, the authors of Huang et al. (2025) compute $\{\hat{p}_{j},\hat{q}_{j}\}_{j=1}^{385}$ for seven LLMs: GPT-3.5-Turbo (gpt-3.5-turbo), GPT-4o (gpt-4o), and GPT-4o-mini (gpt-4o-mini); Claude 3.5 Haiku (claude-3-5-haiku-20241022); Llama 3.3 70B (Llama-3.3-70B-Instruct-Turbo); Mistral 7B (Mistral-7B-Instruct-v0.3); DeepSeek-V3 (DeepSeek-V3), and constructed a baseline random simulator that selects uniformly among the available answer choices. A more detailed exploration into the OpinionQA dataset and the simulation procedure can be found in Huang et al. (2025). We also provide the same procedure onto a Bernoulli setting using the EEDI dataset, details can be found in Appendix E.
We apply our methodology to produce a fidelity profile for each candidate LLM. We use total variation as the discrepancy measure, $L(p,q)=\tfrac{1}{2}\,\|p-q\|_{1}$. In addition, we set the simulation budget $k=100$, the result is presented in Figure 14.
Figure 14: Quantile fidelity profiles $\hat{V}(\alpha)$ across LLMs.
Figure 14 compares models by how tightly their synthetic outcomes track the human distribution across items. We plot $\hat{V}_{\ell}(\alpha)$ against $\alpha$, where lower-flatter curves indicate uniformly small discrepancies, while elbows reveal rare but severe misses. GPT-4o lies lowest across most quantiles, indicating the most reliable alignment. Notably, the simulator curves are steeper than the random benchmark, indicating question-dependent alignment and less uniform discrepancies. This suggests the simulators may require further fine-tuning to achieve more uniform discrepancy levels across this set of questions. Nonetheless, the simulators are, in general, lower than random benchmark, which indicates we have a alignment result than EEDI.